A Bibliography of Publications of Lloyd Nicholas Trefethen

Lloyd Nicholas Trefethen

Cornell University

Upson Hall Ithaca, NY 14853

USA

Tel: +1 607 255 4222 FAX: +1 607 255 4428

E-mail: lnt@cs.cornell.edu (Internet)

25 March 2025 Version 2.94

Abstract

This bibliography records publications of Lloyd Nicholas Trefethen.

Title word cross-reference

$100 [Tre02a]. 10^{7} [MT03]. $59.95 [Tre13b]. A_{α}, log(A) [HHT08]. e^{z} [Tre84a]. ℓ^{p} [Tre84e]. εy′′ − xy′ + y = 0 [Tre20b]. x^{n} [NT18].

-instability [Tre84e].

0 [Böt06]. 0-691-11946-5 [Böt06].

100-Digit [GI02, Tre02c, BLWW04, Tre02a]. 100-Dollar [GI02, Tre02c]. 14th [GW91]. 17th [HWG98]. 1980 [Ano80]. 1981 [CMW82]. 1982 [W^{+}83]. 1983 [CSW83, CMW84, GMSV84]. 1984 [VS84b]. 1985 [MC87]. 1988 [MC90, MB88]. 1995 [KMM96]. 1997 [HWG98].

20.-26 [CMW84]. 2000 [vdVDE^{+}02]. 25th [Tre22c]. 27 [FSR98].

5 [Böt06].

75th [GOR^{+}08].

80th [STW10].

9 [Tre13b]. 978 [Tre13b]. 978-3-0348-0179-9 [Tre13b]. ’98 [ALM99].

= [CMW82, CMW84].

AAA [CT23, DNT24, HT23, NST18, XWT24]. AAA-least [CT23]. absorbing [DT96, TH86a, TH86b, TH88]. accuracy [Tre84a]. Acta [Ise99]. Adaptive [FNTB18]. adaptively [TT06]. Advanced [DMLT84, DL85, LTS85, MN85, TLMD84, TSL85, W^{+}83]. Advances [VS84a, VS84b]. AG [Tre13b]. aktuelle [SL13]. Algebra [TB97b, TB22, Tre90]. Algorithm [NRT92b, NST18, NT20]. Algorithms [AKT14, MC87, MC90, PT09, MC87, MC90]. Analogue [RT11]. analogues [TT15]. Analysis [Ano80, BCC^{+}79, SS03, TA84, Tre84b, ALM99, GW86, GW91, HWG98, LT88, SS93, Tre92a, W^{+}83, ALM99]. analyst [JT98, TJ98]. Analytic [AKT14, PTK11, Tre20d, Tre23a]. angle [HT88]. anniversary [Tre22c]. Apology [Tre22a]. application [Tre80a, Tre81a]. Applications [MB88, SS03]. Applied [KMM96, Tre22a, Tre15b]. Approach [BET03, BET02]. Approximate [ADGP16]. Approximation [CSW83, CMW82, CMW84, FNTB18, GGT13, GT94, NST18, NT21, PTK11, Tre77, Tre90, Tre13a, Tre17a, Tre20a, Cor16, CT23, DNT24, GMSV84, GT82, GT83b, GT84, MC87, MC90, NT18, NT20, PT09, ST08, Tre81b, Tre81c, Tre81d, Tre83a, Tre83b, TG83c, TG83b, TG83d, TG83e, TG84, TG85, Tre86b, TG87, Tre17b, TNW21, VT11, XWT24, CSW83, CSW86, Cor16]. Approximations [ST07, Tre84b, GT83a, TG83a, Tre84a, TWS06]. Approximationstheorie [CMW82, CMW84]. area [BT01]. Arithmetic [AT15]. Army [Ano80]. Arnoldi [BNT21, GT94, TT96a]. Arnoldi/Lánczos [GT94]. ARPACK [WT01b]. aspects [W^{+}83]. association [MB88]. asymptotic [Tre84a]. August [FSR98, W^{+}83]. automatic [DBT08]. Average [TS87, TS90]. Average-Case [TS90].

B [BF81, Tre84b]. Barycentric [BT04b, FNTB18, WTG12, PT09]. Based [AKT14, MC87, MC90, MB88]. Basel [Tre13b]. Bau [Tre22c]. Behavior [Böt06, GT93, TE05, TG97]. behaviour [TV96, Tre99c]. Berlin [FSR98]. best [GT83a, PT09, TG83a, Tre84a]. Bethlehem [VS84b, VS84a]. better [Tre08]. beyond [TTRD92, Tre20e]. bidiagonal [TCE01]. Biennial [HWG98]. bifurcation [Tre23c]. Bill [STW10]. Birkhäuser [Tre13b]. Birkhäuser/Springer [Tre13b]. birthday [GOR^{+}08, STW10]. Block [AT17a]. blocks [Tre84f]. Book [Böt06, Cor16, Tre84b, Tre13b, Tre15b]. bound [JT14]. boundaries [Tre85a, Tre85b]. boundary [DT96, Tre84c, Tre84d, TW86, TH86a, TH86b, TH88]. Bowles [Tre84b]. Braunlage [W^{+}83]. Brief [KW02]. Bromwich [WT07b]. Buffon [WT94]. bugs [CLT11].

Cage [CHT15, Tre16b]. Calculation [ADGP16, TT96a]. California [Ano80]. Carath‘eodory [GT82, GST83, ST08, TG83c, VT11]. card [JT98, TJ98]. cards [TT00b, Tre11c]. Case [TS87, TS90]. CF [HTG90, Tre83b, Tre84f, Tre86b, THG90]. chains [JT98, TJ98]. Challenge [DHN^{+}02, Kel02, Ker02a, Ker02b, MS02, MBvG02, SWRG02, Tre02a, Tre02b, Tre02c, BLWW04, Lau02, GI02, KW02]. Chastened [Tre02a]. Chebfun [DHT14, HT12, HT17, PT10, RT11, TT13, WJMT15, PT09]. chebfuns [PPT10]. chebop [DBT08]. Chebyshev [AT17b, GT82, GT83a, GT83b, GT84, TT06, TT98, Tre77, Tre81b, Tre81c, Tre81d, Tre83a, Tre83b, TG83a, TG83b, TG83d, TG83e, TG84, TG85, TG87]. Chopping [AT17b]. Christoffel [Tre80b, BT03, BF81, DT02, HT90, Tre80a, Tre80c, Tre81a, Tre89a, TD98]. Circularity [Tre83b, Tre81b, Tre81c]. Classical [ET86]. Clenshaw [Tre08, WT07a]. clustering [TNW21]. coefficient [Tre05]. coefficients [Tre23b, WT01a]. College [CSW83, MC87, MC90]. collocation [TT06]. Companion [TT94, Tre15b].

Competitions [SL11]. Complex [SS03, TT00a, Tre77, GT83b, GT84, NT20, SS93, TT99, Tre81b, Tre81c, Tre83b, TG83b, TG83d, TG83e, TG84, Tre13b, W^{+}83]. Computation [Tre80a, Tre99a, WT01b, XWT24, Tre80b, Tre80c]. Computational [MB88, W^{+}83]. computations [EOS85]. Computed [TB06]. Computer [Tre81a, TD98, VS84b, VS84a]. Computers [Ano80]. Computing [AT15, HHT08, ST07, TW06, Tre07, Tre15a, WT01a, PT10, WT07b]. Condition [VT98, LT84]. Conditioned [Tre20b]. conditioning [Tre20d]. conditions [TW86, TH86a, TH86b, TH88]. Conference [Ano80, GW91, HWG98, MC87, MC90, GMSV84, MB88, STW10, vdVDE^{+}02]. Conformal [ET99a, TW86, Tre86c, Tre20c, GT19b, HT08, TA84, Tre86e, Tre93]. Congress [FSR98, KMM96]. Connected [ET99a]. constants [WT01a]. construction [Tre93]. containing [Tre85a]. continuation [Tre20d, Tre23a]. Continuous [BT04a, BET03, TT15, BET02]. continuum [DNT24]. Contour [ST07, HHT08, ST08]. Contours [WT06, WT07b]. Convection [RTP93, RT94]. Convection-Diffusion [RT94]. convergence [JT14, TG85]. Convergent [TW14]. Copper [vdVDE^{+}02]. Corner [GT19c]. Couette [TTR92]. CR [GT94]. Cubature [Tre17a]. current [SL13]. Curtis [Tre08, WT07a]. curve [Tre81b, Tre81c, Tre83b]. cutoff [JT98, TJ98].

Data [MC87, HT23]. decay [ET99b]. December [GMSV84]. deck [TT00b]. dedicated [GOR^{+}08]. definition [Tre92a]. degeneracy [TG85]. derivative [TW86]. design [GST83, TA84]. Determine [GT93, TG97]. Difference [Tre82a, Tre82b, Tre84e, Tre84c, Tre84d, Tre85a, Tre85b]. differential [DBT08, Tre05, VS84b, VS84a]. differentiation [WT88]. Diffusion [RTP93, RT94]. Digit [DHN^{+}02, GI02, KW02, MS02, SWRG02, Tre02b, Tre02c, BLWW04, Kel02, Ker02a, Ker02b, Lau02, Tre02a, Tre11b, Tre13c]. digital [GST83]. dimensional [BT97, TB97a, XWT24]. Dimensions [HT17, TT13, DET87]. direction [TTRD92]. discontinuities [Tre84e]. discrete [RT90]. Discretizations [AT17a]. disk [GT83a, GT83b, Tre81d, Tre83a, TG83a, TG83b, TG83d, Tre84a]. Dispersion [Tre86a, Tre84e]. dissipation [Tre86a]. Do [GT93, TG97]. Dollar [GI02, KW02, MS02, SWRG02, Tre02b, Tre02c, Kel02, Ker02a, Ker02b, Lau02]. Domain [TT00a, TT99]. Domains [ET99a, GT84, TG84]. Dundee [GW91, HWG98]. dynamics [MB88].

E. [BF81]. Editor [Kel02]. eigenmodes [TB06]. Eigenvalue [ET01a, HNT22, Tre88]. Eigenvalues [AT15, HT01, RT92b, TR92b, TH97, RT90, TTRD92, TTRD93, WT88]. Eight [Tre20b]. eigs [WT01b]. Einblicke [SL13]. Einladung [SL13]. Elias [Tre13b]. Elimination [TS87, TS90, Tre85c]. elongated [HT90]. Embree [Böt06]. Engineering [SS03, SS93]. EPSRC [ALM99]. equally [TW91]. Equation [Tre20b, DT96, TW86]. equations [DBT08, HT88, TH86a, TH86b, VS84b, VS84a, Tre84b]. equations-V [VS84a]. equioscillation [Tre84f]. Equispaced [PTK11, HT23]. era [TD98]. Erratum [Tre80b]. error [JT14, Tre81c, Tre83b]. Euler [JT14, JT16]. Evaluating [ST08]. evolution [TTR92]. Exactness [Tre22b]. expansion [Tre23b]. Exploring [TBD18]. Exponential [TNW21, ST08]. Exponentially [TW14, Tre20b]. Extension [BT04a, TT13, WJMT15]. extrapolation [WTG12].

factorizations [TT15]. Faraday [CHT15, Tre16b]. Fast [NRT92a, PTK11, TW06]. February [Ano80]. Fejér [GT82, GST83, ST08, TG83c, VT11, WT07a]. Fibonacci [ET99b]. Field [Ano80]. Fifth [VS84b, VS84a]. filter [GST83]. Filters [AT15]. Finite [Tre82a, Tre82b, Tre84e, Tre84d, Tre85a, Tre85b]. finite-difference [Tre85a, Tre85b]. Florida [GMSV84]. flow [DET87, ET86, MT03, Tre94, TTS99]. flows [TTR92, XWT24]. fluid [EOS85, MB88, TP90]. formula [JT14]. Formulas [Tre22b, HT08, WTG12]. Forschung [SL13]. forty [Tre11c]. Four [CLT11]. Fourier [LT88, Tre84b]. Fourth [KT05]. Fourth-Order [KT05]. free [ET86]. free-streamline [ET86]. fully [RT90]. Function [AKT14, RT11, ST07, BT01]. Functions [BT04a, ET99a, FJT19, GT19c, HHT23, HT24, MC87, PTK11, Tre20c, WJMT15, GT19b, HHT08, JT14, ST08, Tre07, Tre13b, Tre15a, Tre20e]. Fundamentals [SS93, SS03].

Gamma [ST07]. Gateway [ET01b]. Gauss [Tre08]. Gaussian [ADGP16, FJT19, Tre85c, TS87, TS90]. Gene [GOR^{+}08]. General [TT00a, TT99]. Generalizing [ET01a]. geometric [JT14]. German [SL13, Tre13c]. Germany [KMM96, W^{+}83]. GMRES [GT94, NRT92b]. GMRES/CR [GT94]. Golub [GOR^{+}08]. graduate [ALM99]. Green [ET99a]. Gregory [JT16]. grid [TT06]. Group [Tre82a, Tre83c]. Growth [ET99b]. Guide [BCC^{+}79, DHT14, Tre89b, ALM99]. Gustafsson [Tre83c].

Hadamard [TV96]. Hamburg [KMM96]. Harz [W^{+}83]. held [CSW83, GMSV84, KMM96, MC87, MC90, MB88, VS84b, VS84a, W^{+}83]. Helmholtz [GT19a]. Householder [Tre10]. hp [HT24]. hp-Mesh [HT24]. Hundred [DHN^{+}02, KW02, Kel02, Ker02a, Ker02b, Lau02, MS02, SWRG02, Tre02b]. Hundred-Digit [KW02, MS02, SWRG02, Tre02b, Kel02, Ker02a, Ker02b, Lau02]. Hundred-Dollar [KW02, MS02, SWRG02, Tre02b, Kel02, Ker02a, Ker02b, Lau02]. Hybrid [NRT92b]. Hydrodynamic [TTRD93, TTRD92]. Hyperbolic [Tre84b, Tre84c, Tre84d, Tre85b, WT07b]. Hypercube [Tre17a, Tre17b].

ICIAM [KMM96]. Ideal [DET87]. II [MC90, ADGP16]. III [MB88]. Ill [TH88, Tre20b, Tre20d]. Ill-Conditioned [Tre20b]. ill-conditioning [Tre20d]. Ill-posedness [TH88]. IMA [MC87]. IMACS [VS84b, VS84a]. implementation [VT11]. Impossibility [PTK11]. index [Tre11c]. Industrial [KMM96]. Infinity [BET03, BET02]. influence [BF81]. initial [Tre84c, Tre84d]. initial-boundary [Tre84c]. insights [SL13]. Instability [Tre84c, Tre84d, TT87, Tre84e, TTR92]. instead [Tre07, Tre15a]. Institute [MB88, W^{+}83]. instruments [HT01, TH97]. integral [WT07b]. Integrals [ST07, HHT08, ST08]. integrators [ST08]. interfaces [Tre85a]. International [CSW83, FSR98, KMM96, MC90, VS84b, VS84a]. interpolants [JT16]. Interpolation [AT17c, BT04b, GPT11, GMSV84, HT23, TW91, Tre11a, WTG12]. interpretation [Tre83c]. interval [GT83b, TG83b, TG83d, TG83e]. Introduction [Tre00d, Tre00e, Tre02d, Tre13b]. Inverse [Tre16a]. Inversion [WT06]. Invitation [SL11, SL13]. ISBN [Böt06, Tre13b]. Isotropy [Tre17a]. issue [Tre86e]. iteration [LT88, TT96a]. Iterations [DTT98, NRT92a, TDTxx]. IV [CSW83].

J [Tre84b]. Januar [CMW82]. January [CSW83, CMW82]. jet [DET87]. July [KMM96, MC87, MC90, W^{+}83]. June [GW91, HWG98, VS84b, VS84a].

kind [PT10]. Kingdom [GMSV84]. kink [WT07a]. Kreiss [LT84, TT99, TT00a, Tre83c, WT94].

L. [Cor16]. Lagrange [BT04b]. Lánczos [GT94]. Laplace [CT23, GT19a, GT19c, TW86, Tre18, WT06]. Large [EOS85, WT01b, Tre23b]. Large-Scale [WT01b, EOS85]. law [Tre23c]. Lax [RT90, Tre88]. Lax-stability [RT90, Tre88]. least [CT23, GPT11]. least-squares [GPT11]. lecture [ALM99]. Lehigh [VS84b, VS84a]. Letter [Kel02]. Library [BCC^{+}79]. Lightning [BT22, XWT24]. Linear [NRT92b, TB97b, Tre97, TB22, BDT95, Tre90, TTR92, Tre96]. Linearized [MT03]. lines [RT92a, TR92a]. Lloyd [Böt06]. localization [TCE01]. Log [NT21]. looks [JT98, TJ98]. Low [BT97, TB97a]. Low-dimensional [BT97, TB97a]. Lyapunov [WT01a].

M [CSW83]. Maclaurin [JT14, JT16]. many [TT00b]. Mapping [BT03, ET99a, Tre86c, Tre20c, DT02, TA84, TW86, Tre86e, Tre89a, TD98]. maps [GT19b, HT08, Tre93]. March [CMW84, MB88]. Mark [Böt06]. Markov [JT98, TJ98]. März [CMW84]. Mathematician [Tre22a]. Mathematicians [FSR98]. Mathematics [CHT15, KMM96, MB88, SL11, BF81, SS93, SL13, Tre11c, Tre15b]. Mathematik [SL13]. MATLAB [BT04a, Tre86b, TT^{+}96b, TMC^{+}99, Tre00a]. Matrices [AT15, BET03, Böt06, TT94, TE05, VT98, BET02, RT92b, TR92b, Tre92b, Tre99c, TCE01, TC04, WT88, WT02]. Matrix [DTT98, GT93, GT94, LT84, NRT92a, TT98, TT00a, TV96, HHT08, ST08, TT99, TT15, TG97, TDTxx, WT94]. mechanics [EOS85, TP90]. Mesh [HT24]. Method [BT03, BT05, GT82, GST83, RT92a, TT06, Tre83b, TG83c, TR92a, VT11]. Methoden [CMW82, CMW84]. Methods [CMW82, CMW84, TT87, Tre00a, MB88, RT90, Tre88, VS84b, VS84a]. migration [TH88]. Military [MC87, MC90]. Minimax [FNTB18, NT20]. Misjudged [Tre02a]. model [BDT95]. models [BT97, Tre84c, Tre84d, Tre85a, Tre85b, TB97a]. modified [HT90]. Moffett [Ano80]. Moore [Tre23c]. Morton [STW10]. mostly [BDT95]. Mountain [vdVDE^{+}02]. MR3380576 [Tre15b]. MultiMATLAB [TT^{+}96b, TMC^{+}99]. Multiple [TMC^{+}99, TT^{+}96b]. Multiply [ET99a]. Multipole [BT03]. Multiscale [HT24]. Multivariate [Tre17b]. Müntz [Tre23b]. musical [HT01, TH97]. mysteries [Tre85c]. myths [Tre11a].

N [Böt06, Cor16]. NAPLUG [BCC^{+}79]. NATO [W^{+}83]. Navier [TTR92]. Near [Tre81b, Tre81c]. Near-circularity [Tre81b, Tre81c]. needle [WT94]. node [TNW21]. non [Tre99c]. non-normal [Tre99c]. Nonhermitian [Tre06]. Nonnormal [Böt06, TE05]. Nonsymmetric [NRT92a, NRT92b]. Nonuniqueness [GT83a, TG83a]. normal [Tre99c]. notes [ALM99, Tre11c]. number [MT03]. Numbers [VT98, Tre07, Tre15a]. numerica [Ise99]. Numerical [Ano80, AKT14, BT01, BCC^{+}79, CMW82, CMW84, GW86, GW91, Tre80b, Tre80c, Tre84b, Tre86c, Tre93, TB97b, Tre20c, TB22, Tre23a, ALM99, HT12, JT98, MB88, PT10, Tre86e, Tre90, Tre92a, TJ98, HWG98, MB88]. numerically [Tre07, Tre15a]. Numerische [CMW82, CMW84].

Oberwolfach [CMW82, CMW84]. oblique [TW86]. obstacle [ET86]. occasion [GOR^{+}08, Tre22c]. ODEs [FJT19, HT93, TBD18]. omitted [BT01]. one [HT88, Tre85b, TH86a, TH86b]. one-way [HT88, TH86a, TH86b]. Operator [RTP93, RT94, TTR92]. Operators [AT17a, BT04a, BET03, Böt06, Tre97, TE05, BET02, Tre96, Tre99c, Tre05]. Optimizing [WT06]. Order [KT05, WT88]. organized [MB88]. other [JT98, TJ98]. Oxford [Böt06, MB88].

packet [TC04, Tre05]. Padé [GGT13, Tre84f, TG85, TG87]. pages [Böt06]. Parabolic [WT07b]. partial [VS84b, VS84a]. Particular [BT05]. PDE [NT21]. PDEs [KT05, TNW21]. Pennsylvania [VS84b, VS84a]. people [Tre11c]. Periodic [WJMT15, JT14]. Perspectives [Tre20b]. Perturbed [AT17c]. phase [Tre13b]. Phenomenon [TT87, JT98, TJ98, WT07a]. physical [BF81]. Piecewise [BET02, BET03, PPT10]. Piecewise-smooth [PPT10]. pipe [MT03, Tre94, TTS99]. Planar [NT21, TB06]. Plane [Tre77, TTR92]. Points [AT17c, TT06, TW91]. Poiseuille [TTR92, Tre94, TTS99]. polygon [TW86]. polygonal [ET86, TA84]. Polygons [BT03]. polynomial [GT82, PT09, TW91, Tre11a, Tre17b]. Polynomials [TT94, TT98, Tre77]. portraits [Tre13b]. posedness [TH86a, TH86b, TH88]. Potential [DTT98, TDTxx]. pp [Tre13b]. Practice [Tre13a, Tre20a, Cor16]. Preface [GOR^{+}08, Tre86d]. Press [Böt06]. Princeton [Böt06, Tre15b]. problem [BT01, WT94]. Probleme [Tre13c]. Problems [DMLT84, DL85, GT19c, GT94, LTS85, MN85, TLMD84, TSL85, CT23, HNT22, Lau02, Tre84c, Tre84d, Tre11b, Tre13c, Tre18]. proceedings [CSW83, GMSV84, GW91, HWG98, KMM96, MC87, MC90, MB88, VS84b, VS84a, W^{+}83, Ano80, FSR98]. Processes [FJT19]. Processors [TMC^{+}99, TT^{+}96b]. products [TV96]. Program [BCC^{+}79]. programs [Tre86b]. Propagation [Tre82b]. pseudo [RT90, RT92b, TR92b]. pseudo-eigenvalues [RT90, RT92b, TR92b]. pseudomodes [TC04, Tre05]. Pseudospectra [Böt06, DT96, ET01a, ET01b, GT93, RT94, TT94, TTR92, Tre92b, Tre96, Tre97, Tre99b, TE05, WT01b, WT02, TT96a, Tre94, TG97, Tre99a, Tre99c, TTS99, TCE01, Tre06]. Pseudospectrum [RTP93]. Pseudozeros [TT94].

Quadrature

[ADGP16, AT17c, Tre22b, HT08, HT12, Tre08, Tre11a, TNW21, WT07a]. quadratures [TWS06]. Quantifying [Tre20d]. quasimatrix [Tre10]. questions [TP90].

R [Tre84b]. Random

[FJT19, VT98, ET99b, TCE01, WT01a].

randomize [TT00b]. Rational

[AT15, FNTB18, GT19c, HHT23, NST18, NT18, ST07, Tre81d, Tre20c, Tre20e, CT23, DNT24, GPT11, GT19b, GT83a, GT84, NT20, TT06, TG83c, TG83a, TG83e, Tre84a, TG84, TG85, TWS06, TNW21, VT11, XWT24, GMSV84]. Reading [MB88]. Real [AT15, GT82, GT83b, GT84, TG83b, TG83d, TG83e, TG84, NT20, TG83c]. Reciprocal [NT21]. Reciprocal-Log [NT21]. recollections [Tre22c]. rectangle [CLT11]. Rectangular [HNT22, WT02]. recurrences [WT01a]. recursive [GST83]. reducibility [TV96]. Refinement [HT24]. regions [HT90, RT90, TB06]. related [HHT08]. Remez [PT09]. Representation [GT19b]. Representations [FNTB18]. Research [SL11, SL13]. resistors [TA84].

Resolution [HHT23, HT24]. resolvent [LT84]. results [TW91]. Review [Böt06, Tre84b, Tre99d, Tre99e, Tre99f, Tre99g, Tre00b, Tre00c, Tre01a, Tre01b, Tre01c, Tre01d, Tre02e, Tre02f, Tre02g, Tre02h, Tre13b, Cor16, Tre00d, Tre00e, Tre02d, Tre15b]. Reviving [BT05]. Reynolds [MT03]. Robust [GPT11, GGT13, VT11]. Roots [AKT14]. Royal [MC87, MC90]. Rule [TW06, TW14, JT14].

Samples [PTK11]. Scale [WT01b, EOS85]. Schemes [Tre82a, Tre82b, Tre84e]. School [ALM99]. Schwarz [Tre80b, BT03, DT02, HT90, Tre80a, Tre80c, Tre81a, Tre89a, TD98]. Science [MC87, MC90, SS03, SS93]. sciences [BF81]. Scientific [TW06]. SCPACK [Tre89b]. Second [MC90, WT88]. second-order [WT88]. sequences [ET99b]. Series [AT17b, Tre18]. sharpness [Tre83b]. Shrivenham [MC87, MC90]. shuffles [TT00b]. shuffling [JT98, TJ98]. SIAM [BLWW04, GI02, KW02, Ker02b, MBvG02, Tre02c]. Sides [BT03]. Sigmoid [HT24]. Sinc [RT11]. Singularities [GT19c, HHT23, HT24, TNW21]. Six [DTT98, Tre11a, TDTxx]. slow [BET02, BET03]. Smooth [FJT19, PPT10, WT01a]. softcover [Tre13b]. Solution [Ker02a, Ker02b, BT01, CT23, DBT08, TW86, Tre18]. Solutions [BT05, DMLT84, DL85, DHN^{+}02, GI02, KW02, Lau02, LTS85, MN85, MS02, TLMD84, TSL85]. solve [SWRG02]. Solver [BT22]. Solvers [NT21, GT19a]. Solving [GT19c]. Some [TP90, Tre22c]. SOR [LT88]. spaced [TW91]. Special [Tre86e, GOR^{+}08]. Spectacularly [Tre23b]. Spectra [Böt06, Tre94, Tre99b, Tre99c, TTS99, TCE01, TE05]. Spectral [AT17a, TT87, Tre00a, RT90, TT06, Tre88, WT88]. Sponsor [Tre02a]. Springer [Tre13b].

Square [Tre84f]. squares [CT23, GPT11]. Stability [RT92a, Tre82b, Tre85a, Tre85b, TS87, TS90, TR92a, WTG12, RT90, Tre83c, Tre86a, Tre88, TTRD92, TTRD93, RT90]. Stable [PTK11, TG87]. States [GMSV84]. Station [CSW83]. Stepping [KT05]. Steps [DTT98, TDTxx]. Stiff [KT05]. Stiffness [HT93]. Stokes [BT22, TTR92, XWT24]. streamline [ET86]. student [ALM99]. Study [W^{+}83]. subcritical [BT97, TB97a]. Summer [ALM99]. Sums [ADGP16]. Sundström [Tre83c]. Surprises [Tre16b]. Survey [Tre99d, Tre99e, Tre99f, Tre99g, Tre00d, Tre00e, Tre00b, Tre00c, Tre01a, Tre01b, Tre01c, Tre01d, Tre02d, Tre02e, Tre02f, Tre02g, Tre02h]. SVD [GGT13]. Symmetric [AT15]. Symposium [CSW83, VS84b, VS84a]. system [DBT08, PT09]. Systems [NRT92b, Tre06].

table [HTG90, THG90]. tables [Tre84f]. Tagung [CMW82, CMW84]. Talbot [TWS06, WT06]. Tampa [GMSV84]. Ten [Tre11b, Tre13c]. Ten-digit [Tre13c]. Texas [CSW83]. Theorem [TT00a, TT99, Tre23b, WT94, LT84]. Theorems [ET01a]. Theory [CMW82, CMW84, DTT98, Tre13a, BT01, CSW83, CSW86, Cor16, Tre83c, Tre90, TDTxx, Tre20a]. Third [KMM96]. Thousands [BT03]. Three [HT17, Tre85c]. Time [KT05]. Time-Stepping [KT05]. Toeplitz [BET02, BET03, RT92b, TR92b, TC04]. Transform [WT06]. transformation [HT90, Tre80b, Tre80c, Tre81a]. transformations [Tre80a]. transformed [TT06]. transition [BDT95, BT97, TB97a]. Trapezoid [TW06]. Trapezoidal [TW14, JT14]. Trefethen [Böt06, Cor16, Lau02, Tre11c, Tre22c]. Triangular [VT98]. triangularization [Tre10]. Trigonometric [AT17c]. turbulence [BDT95, BT97, TB97a]. twisted [TC04]. Two [TT13, TW91, DET87, Tre85a, Tre85b, XWT24]. two-dimensional [XWT24]. Type [ADGP16].

unanswered [TP90]. unifying [JT14]. unit [GT83a, GT83b, Tre81d, Tre83a, TG83a, TG83b, TG83d]. unitary [TV96]. United [GMSV84]. Unity [AKT14]. univalent [BT01]. Universities [MB88]. University [Böt06, CSW83, VS84b, VS84a]. USA [VS84b, VS84a]. User [BCC^{+}79, Tre89b]. Using [ST07, WT01b].

V [VS84b, CSW86, VS84a]. value [Tre84c, Tre84d]. Values [AKT14]. Vandermonde [BNT21]. variable [Tre05]. Velocity [Tre82a, Tre83c]. via [ET99a, FNTB18, GGT13, GT19c, RT90, ST08, XWT24]. Vichnevetsky [Tre84b]. VIII [ALM99]. Visual [Tre13b]. volume [GOR^{+}08]. vs [GT84, TG83e, TG84, Tre88].

Walsh [Tre84f]. Wave [Tre82b, TC04, Tre05, DT96, HT88, TH86a, TH86b]. way [HT88, TH86a, TH86b]. Wegert [Tre13b]. Well [TH86a, TH86b]. Well-posedness [TH86a, TH86b]. Wide [HT88]. Wide-angle [HT88]. without [TTRD93]. words [Tre11c]. work [BF81]. Workshop [CMW82, CMW84].

xiv [Tre13b]. xvii [Böt06].

years [Tre11c]. Yogiisms [Tre16a].

Zehnstellige [Tre13c].

References

Area:2016:ACS

[ADGP16] Iván Area, Dimitar K. Dimitrov, Eduardo Godoy, and Vanessa G.

Paschoa. Approximate calculation of sums II: Gaussian type quadrature. SIAM Journal on Numerical Analysis, 54(4):2210– 2227, ???? 2016. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Austin:2014:NAB

[AKT14] Anthony P. Austin, Peter Kravanja, and Lloyd N. Trefethen. Numerical algorithms based on analytic function values at roots of unity. SIAM Journal on Numerical Analysis, 52(4):1795– 1821, ???? 2014. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:1999:GSG

[ALM99] M. Ainsworth, Jeremy Levesley, and Marco Marletta, editors. The graduate student’s guide to numerical analysis ’98: lecture notes from the VIII EPSRC Summer School in Numerical Analysis, volume 26 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65752-5. ISSN 0179-3632. LCCN QA297 .E57 1998.

Anonymous:1980:PAN

[Ano80] Anonymous, editor. Proceedings of the 1980 Army Numerical Analysis and Computers Conference, Moffett Field, California, February 20–21, 1980, ARO Rep. 80. U.S. Army Research Office, Research Triangle Park, NC, USA, 1980. ISBN ???? LCCN ???? Six microfiches.

Austin:2015:CER

[AT15] Anthony P. Austin and Lloyd N.

Trefethen. Computing eigenvalues of real symmetric matrices with rational filters in real arithmetic. SIAM Journal on Scientific Computing, 37(3): A1365–A1387, ???? 2015. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Aurentz:2017:BOS

[AT17a] Jared L. Aurentz and Lloyd N.

Trefethen. Block operators and spectral discretizations. SIAM Review, 59(2):423–446, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aurentz:2017:CCS

[AT17b] Jared L. Aurentz and Lloyd N.

Trefethen. Chopping a Chebyshev series. ACM Transactions on Mathematical Software, 43(4):33:1–33:21, March 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Austin:2017:TIQ

[AT17c] Anthony P. Austin and Lloyd N.

Trefethen. Trigonometric interpolation and quadrature in perturbed points. SIAM Journal on Numerical Analysis, 55(5):2113– 2122, ???? 2017. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bolstad:1979:NAP

[BCC^{+}79] J. H. Bolstad, T. F. Chan, W. M. Coughran, Jr., W. D. Gropp, E. H. Grosse, M. T. Heath, R. J. LeVeque, F. T. Luk, S. G. Nash, and L. N. Trefethen. Numerical analysis program library user’s guide (NAPLUG). User Note 82, SLAC Computing Services, Stanford, CA, USA, 1979. First issued in 1976 by Chan, Coughran, Heath, and Luk.

Baggett:1995:MLM

[BDT95] Jeffrey S. Baggett, Tobin A. Driscoll, and Lloyd N. Trefethen. A mostly linear model of transition to turbulence. Physics of Fluids, 7(4):833–838, ???? 1995. CODEN PHFLE6. ISSN 1070-6631.

Bottcher:2002:PCT

[BET02] Albrecht Böttcher, Mark Embree, and Lloyd N. Trefethen. Piecewise continuous Toeplitz matrices and operators: slow approach to infinity. SIAM Journal on Matrix Analysis and Applications, 24(2):484–489, 2002. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Bottcher:2003:PCT

[BET03] Albrecht Böttcher, Mark Embree, and Lloyd N. Trefethen. Piecewise continuous Toeplitz matrices and operators: Slow approach to infinity. SIAM Journal on Matrix Analysis and Applications, 24(2):484– 489, April 2003. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/37697.

Butzer:1981:BCI

[BF81] Paul Leo Butzer and F. (Franziska)

Fehér, editors. E. B. Christoffel, the influence of his work on mathematics and the physical sciences. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1981. ISBN 3-7643-1162-2. LCCN QA7 I57 1981; QA 7 .I57 1981; QA76 .I556 1981. URL https:

//link.springer.com/book/
10.1007/978-3-0348-5452-8.

Bornemann:2004:SDC

[BLWW04] Folkmar Bornemann, Dirk Laurie, Stan Wagon, and Jörg Waldvogel. The SIAM 100-digit challenge. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2004. ISBN 0-89871-561-X (paperback). xii + 306 pp. LCCN QA297 .S4782 2004. US$57.00. URL http://www-

m8.mathematik.tu-muenchen.
de/m3/ftp/Bornemann/pdf/story.

pdf. A study in high-accuracy numerical computing, with a foreword by David H. Bailey. See [Tre02b].

Brubeck:2021:VA

[BNT21] Pablo D. Brubeck, Yuji Nakatsukasa, and Lloyd N. Trefethen. Vandermonde with Arnoldi. SIAM Review, 63(2): 405–415, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bottcher:2006:SPB

[Böt06] Albrecht Böttcher. Book review: Spectra and Pseudospectra: The Behavior of Nonnormal Matrices and Operators by Lloyd N. Trefethen and Mark Embree, Princeton University Press, Princeton and Oxford, 2005, xvii + 606 pages, ISBN 0-691-11946-5. Linear Algebra and its Applications, 416(2–3): 1098–1101, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baggett:1997:LDM

[BT97] Jeffrey S. Baggett and Lloyd N. Trefethen. Low-dimensional models of subcritical transition to turbulence. Physics of Fluids, 9(4):1043–1053, 1997. CODEN PHFLE6. ISSN 1070-6631.

Banjai:2001:NSO

[BT01] Lehel Banjai and Lloyd N. Trefethen. Numerical solution of the omitted area problem of univalent function theory. Computational Methods and Function Theory, 1(1):259–273, 2001. CODEN ???? ISSN 1617-9447.

Banjai:2003:MMS

[BT03] Lehel Banjai and L. N. Trefethen. A multipole method for Schwarz–Christoffel mapping of polygons with thousands of sides. SIAM Journal on Scientific Computing, 25 (3):1042–1065, May 2003. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/41167.

Battles:2004:EMC

[BT04a] Zachary Battles and Lloyd N.

Trefethen. An extension of MATLAB to continuous functions and operators. SIAM Journal on Scientific Computing, 25(5):1743–1770, September 2004. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43012.

Berrut:2004:BLI

[BT04b] Jean-Paul Berrut and Lloyd N.

Trefethen. Barycentric Lagrange interpolation. SIAM Review, 46(3):501–517, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41771.

Betcke:2005:RMP

[BT05] Timo Betcke and Lloyd N. Trefethen. Reviving the method of particular solutions. SIAM Review, 47(3):469–491, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43733.

Brubeck:2022:LSS

[BT22] Pablo D. Brubeck and Lloyd N. Trefethen. Lightning Stokes solver. SIAM Journal on Scientific Computing, 44(3):A1205– A1226, ???? 2022. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (elec-

tronic). URL https://epubs.
siam.org/doi//doi/10.1137/
21M1408579.

Chapman:2015:MFC

[CHT15] S. Jonathan Chapman, David P. Hewett, and Lloyd N. Trefethen. Mathematics of the Faraday cage. SIAM Review, 57 (3):398–417, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:2011:FBR

[CLT11] S. J. Chapman, James Lottes, and Lloyd N. Trefethen. Four bugs on a rectangle. Proceedings of the Royal Society A: Mathematical, Physical, & Engineering Sciences, 467(2127): 881–896, March 8, 2011. CO-DEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Collatz:1982:WNM

[CMW82] L. (Lothar) Collatz, Günther Meinardus, and H. (Helmut) Werner, editors. Workshop on Numerical Methods of Approximation Theory, Oberwolfach, January 18–24, 1981 = Tagung über Numerische Methoden der Approximationstheorie, Oberwolfach, 18.–24. Januar 1981, volume 59 of Numerical methods of approximation theory; International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1982. ISBN 3-7643-1304-8. LCCN QA221 .T129n 1981; QA297.5 .N85.

Collatz:1984:WNM

[CMW84] L. (Lothar) Collatz, Günther Meinardus, and H. (Helmut) Werner, editors. Workshop on Numerical Methods of Approximation Theory, Oberwolfach, March 20–26, 1983 = Tagung über Numerische Methoden der Approximationstheorie, Oberwolfach, 20.-26. März 1983, volume 67 of Numerical methods of approximation theory; International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1984. ISBN 3-7643-1580-6. LCCN QA221 .T129n 1983.

Corless:2016:BRT

[Cor16] Robert M. Corless. Book review of: L. N. Trefethen, Approximation theory and approximation practice. SIAM Review, 58(1):159–163, 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1983:ATI

[CSW83] C. K. Chui, Larry L. Schumaker, and J. D. Ward, editors. Approximation theory IV: proceedings of the International Symposium on Approximation Theory held at Texas A and M University, College Station, Texas, on January 10–14, 1983. Academic Press, New York, NY, USA, 1983. ISBN 0-12-174580-5. LCCN QA297.5 .I54 1983.

Chui:1986:ATV

[CSW86] C. K. Chui, Larry L. Schumaker, and J. D. Ward, editors. Approximation theory V. Academic Press, New York, NY, USA, 1986. ISBN 0-12-174581-3. LCCN QA221 .S92 1986; QA221.S92. Proceedings of the Fifth International Symposium on Approximation Theory, held at Texas A and M University on January 13–17, 1986.

Costa:2023:ALS

[CT23] Stefano Costa and Lloyd N. Trefethen. AAA-least squares rational approximation and solution of Laplace problems. In European Congress of Mathematics, pages 511–534. EMS Press, Berlin, Germany, 2023. ISBN 3-9854705-1-0; 3-9854755-1-2.

Driscoll:2008:CSA

[DBT08] Tobin A. Driscoll, Folkmar Bornemann, and Lloyd N. Trefethen. The chebop system for automatic solution of differential equations. BIT, 48(4): 701–723, 2008. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Dias:1987:IJF

[DET87] Frédéric Dias, Alan R. Elcrat, and Lloyd N. Trefethen. Ideal jet flow in two dimensions. J. Fluid Mech., 185:275–288, 1987. CO-DEN JFLSA7. ISSN 0022-1120 (print), 1469-7645 (electronic).

Dussaud:2002:HDC

[DHN^{+}02] Eric Dussaud, Chris Husband, Hoang Nguyen, Dan Reynolds, and Christiaan Stolk. Hundred digit challenge solutions. Technical report, Department of Computational and Applied Mathematics, Rice University, Houston, TX, USA, May 16, 2002. 21 pp. URL

http://www.caam.rice.edu/
caam/trs/2002/TR02-06.pdf;
 http://www.caam.rice.edu/
caam/trs/2002/TR02-06.ps;
 http://www.caam.rice.edu/
caam/trs/2002_abstracts.html#

TR02-06. See [Tre02b, Tre02a, Tre02c].

Driscoll:2014:CG

[DHT14] Tobin A. Driscoll, Nicholas Hale, and Lloyd N. Trefethen, editors. Chebfun Guide. Pafnuty Publications, Oxford, UK, 2014. x + 202 (various paging)

pp. URL http://www.chebfun.
org/docs/guide/.

Duffy:1985:PSS

[DL85] L. Richard Duffy and Detlef Laugwitz. Problems and solutions: Solutions of advanced problems: 6460. Amer. Math. Monthly, 92(9):671, November 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DMLT84].

Duffy:1984:PSA

[DMLT84] L. Richard Duffy, L. E. Mattics, Randall J. LeVeque, and Lloyd N. Trefethen. Problems and solutions: Advanced problems: 6460–6462. Amer. Math. Monthly, 91(6):371, June/July 1984. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DL85, LTS85, MN85].

Driscoll:2024:ARA

[DNT24] Tobin A. Driscoll, Yuji Nakatsukasa, and Lloyd N. Trefethen. AAA rational approximation on a continuum. SIAM Journal on Scientific Computing, 46(2): A929–A952, 2024. CODEN SJOCE3. ISSN 1064-8275,1095-

                                        7197. URL https://doi.org/
                                        10.1137/23M1570508.

Driscoll:1996:PWE

[DT96] Tobin A. Driscoll and Lloyd N. Trefethen. Pseudospectra for the wave equation with an absorbing boundary. Journal of Computational and Applied Mathematics, 69(1):125– 142, April 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0377042795000216.

Driscoll:2002:SCM

[DT02] Tobin A. Driscoll and Lloyd N. Trefethen. Schwarz–Christoffel mapping, volume 8 of Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-80726-3. xvi + 132 pp. LCCN QA360 .D75 2002.

Driscoll:1998:PTM

[DTT98] Tobin A. Driscoll, Kim-Chuan C.

Toh, and Lloyd N. Trefethen. From potential theory to matrix iterations in six steps. SIAM Review, 40(3):547–578, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30558.

Engquist:1985:LSC

[EOS85] Björn Engquist, Stanley Osher, and Richard Somerville, editors. Large-scale computations in fluid mechanics, volume 22 of Lectures in applied mathematics. Amer. Math. Soc., Providence, RI, USA, 1985. ISBN 0-8218-1122-3. LCCN QA901 .L37 1985.

Elcrat:1986:CFS

[ET86] Alan R. Elcrat and Lloyd N.

Trefethen. Classical freestreamline flow over a polygonal obstacle. Journal of Computational and Applied Mathematics, 14(1–2):251–265, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0377042786901421.

Embree:1999:GFM

[ET99a] Mark Embree and Lloyd N. Trefethen. Green’s functions for multiply connected domains via conformal mapping. SIAM Review, 41(4):745–761, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34927.

Embree:1999:GDR

[ET99b] Mark Embree and Lloyd N. Trefethen. Growth and decay of random Fibonacci sequences. Proceedings of the Royal Society A: Mathematical, Physical, & Engineering Sciences, 455 (1987):2471–2485, July 8, 1999. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Embree:2001:GET

[ET01a] Mark Embree and Lloyd N. Trefethen. Generalizing eigenvalue theorems to pseudospectra theorems. SIAM Journal on Scientific Computing, 23(2): 583–590, March 2001. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/37301.

Embree:2001:PG

[ET01b] Mark Embree and Lloyd N. Trefethen. Pseudospectra gateway. Web site, 2001.

                                        URL http://www.comlab.ox.
                                        ac.uk/pseudospectra.

Filip:2019:SRF

[FJT19] Silviu Filip, Aurya Javeed, and Lloyd N. Trefethen. Smooth random functions, random ODEs, and Gaussian processes. SIAM Review, 61(1):185–205, ???? 2019. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Filip:2018:RMA

[FNTB18] Silviu-Ioan Filip, Yuji Nakatsukasa, Lloyd N. Trefethen, and Bernhard Beckermann. Rational minimax approximation via adaptive barycentric representations. SIAM Journal on Scientific Computing, 40(4):A2427– A2455, ???? 2018. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Fischer:1998:PIC

[FSR98] Gerd Fischer, Hannes Stoppel, and Ulf Rehmann, editors. Proceedings of the International Congress of Mathematicians, Berlin 1998, August 18–27, Documenta Mathematica. Deutscher Mathematiker-Vereinigung, Bielefeld, Germany, 1998. ISSN 1431-0635. LCCN QA1 .I61 1998.

Gonnet:2013:RPA

[GGT13] Pedro Gonnet, Stefan Güttel, and Lloyd N. Trefethen. Robust Padé approximation via SVD. SIAM Review, 55(1): 101–117, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gonnet:2002:SSD

[GI02] Gaston Gonnet and Robert Israel. Solutions to the SIAM 100-Dollar, 100-Digit Challenge. World-Wide Web document.,

2002. URL http://www.math.
ubc.ca/~israel/challenge/
challenge.html. See [Tre02b,

Tre02a, Tre02c].

Graves-Morris:1984:RAI

[GMSV84] P. R. Graves-Morris, E. B. Saff, and Richard S. Varga, editors. Rational approximation and interpolation: proceedings of the United Kingdom– United States conference held at Tampa, Florida, December 12– 16, 1983, volume 1105 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-13899-4 (U.S.). LCCN QA3 .L28 no. 1105; QA297. DM72.00.

Gutknecht:2007:PSV

[GOR^{+}08] Martin Gutknecht, Michael Overton, Lothar Reichel, Daniel B. Szyld, Nick Trefethen, Paul Van Dooren, and Andy Wathen. Preface: Special volume dedicated to Gene Golub on the occasion of his 75th birthday. Electronic Transactions on Numerical Analysis, 28:vii, 2007/2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). Special volume for Gene Golub.

Gonnet:2011:RRI

[GPT11] Pedro Gonnet, Ricardo Pachón, and Lloyd N. Trefethen. Robust rational interpolation and least-squares. Electronic Transactions on Numerical Analysis, 38:146–167, 2011. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic).

Gutknecht:1983:CFM

[GST83] Martin H. Gutknecht, Julius O.

Smith, and Lloyd N. Trefethen. The Carathéodory– Fejér method for recursive digital filter design. IEEE Trans. Acoust., Speech, and Signal Process., 31(6):1417–1426, ???? 1983. CODEN IETABA. ISSN 0096-3518. URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1164242.

Gutknecht:1982:RPC

[GT82] Martin H. Gutknecht and Lloyd N. Trefethen. Real polynomial Chebyshev approximation by the Carath‘eodory–Fejér method. SIAM Journal on Numerical Analysis, 19(2):358–371, April 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gutknecht:1983:NBR

[GT83a] Martin H. Gutknecht and Lloyd N. Trefethen. Nonuniqueness of best rational Chebyshev approximations on the unit disk. Journal of Approximation Theory, 39(3):275–288, 1983. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Gutknecht:1983:RCC

[GT83b] Martin H. Gutknecht and Lloyd N. Trefethen. Real and complex Chebyshev approximation on the unit disk and interval. Bulletin of the American Mathematical Society (new series), 8(3):455–458, 1983. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic). URL http:/

/projecteuclid.org/euclid.
bams/1183550890.

Gutknecht:1984:RVC

[GT84] Martin H. Gutknecht and Lloyd N. Trefethen. Real vs. complex rational Chebyshev approximation on complex domains. In Collatz et al. [CMW84], pages 87–97. ISBN 3-7643-1580-6. LCCN QA221 .T129n 1983.

Greenbaum:1993:DPM

[GT93] Anne Greenbaum and Lloyd N. Trefethen. Do the pseudospectra of a matrix determine its behavior? Technical Report TR 93-1371, Department of Computer Science, Cornell University, Ithaca, New York, August 1993. ???? pp.

Greenbaum:1994:GCA

[GT94] Anne Greenbaum and Lloyd N. Trefethen. GMRES/CR and Arnoldi/Lánczos as matrix approximation problems. SIAM Journal on Scientific Computing, 15(2):359–368, March 1994. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). Iterative methods in numerical linear algebra (Copper Mountain Resort, CO, 1992).

Gopal:2019:NLH

[GT19a] Abinand Gopal and Lloyd N.

Trefethen. New Laplace and Helmholtz solvers. Proceedings of the National Academy of Sciences of the United States of America, 116(21):10223–10225, 2019. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Gopal:2019:RCM

[GT19b] Abinand Gopal and Lloyd N.

Trefethen. Representation of conformal maps by rational functions. Num. Math, 142 (2):359–382, 2019. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gopal:2019:SLP

[GT19c] Abinand Gopal and Lloyd N.

Trefethen. Solving Laplace problems with corner singularities via rational functions. SIAM Journal on Numerical Analysis, 57(5):2074–2094, ???? 2019. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Griffiths:1986:NA

[GW86] D. F. (David Francis) Griffiths and G. A. (G. Alistair) Watson, editors. Numerical analysis, volume 140 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, 1986. ISBN 0-470-20669-1, 0-582-98897-7. LCCN QA297 .N8251 1986. The 11th Dundee Biennial Conference on Numerical Analysis, held at the University of Dundee on 25–28 June 1985.

Griffiths:1991:NAP

[GW91] D. F. Griffiths and G. A. Watson, editors. Numerical analysis, 1991: proceedings of the 14th Dundee Conference, June 1991, volume 260 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, 1991. ISBN 0-582-08908-5. LCCN QA297 .D85 1991.

Hale:2008:CRM

[HHT08] Nicholas Hale, Nicholas J. Higham, and Lloyd N. Trefethen. Computing A_{α}, log(A), and related matrix functions by contour integrals. SIAM Journal on Numerical Analysis, 46(5): 2505–2523, 2008. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Herremans:2023:RSR

[HHT23] Astrid Herremans, Daan Huybrechs, and Lloyd N. Trefethen. Resolution of singularities by rational functions. SIAM Journal on Numerical Analysis, 61 (6):2580–2600, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hashemi:2022:REP

[HNT22] Behnam Hashemi, Yuji Nakatsukasa, and Lloyd N. Trefethen. Rectangular eigenvalue problems. Advances in Computational Mathematics, 48(6):1– 16, 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Halpern:1988:WAO

[HT88] Laurence Halpern and Lloyd N.

Trefethen. Wide-angle one-way wave equations. Journal of the Acoustical Society of America, 84(4):1397–1404, October 1988. CODEN JASMAN. ISSN 0001-4966.

Howell:1990:MSC

[HT90] Louis H. Howell and Lloyd N.

Trefethen. A modified Schwarz– Christoffel transformation for elongated regions. SIAM Journal on Scientific and Statistical Computing, 11(5):928–949, September 1990. CODEN SI-JCD4. ISSN 0196-5204.

Higham:1993:SO

[HT93] Desmond J. Higham and Lloyd N.

Trefethen. Stiffness of ODEs. BIT, 33(2):285–303, 1993. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

mai.liu.se/BIT/contents/bit33.
html.

Howle:2001:EMI

[HT01] V. E. Howle and Lloyd N. Trefethen. Eigenvalues and musical instruments. Journal of Computational and Applied Mathematics, 135(1):23–40, October 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0377042700005604.

Hale:2008:NQF

[HT08] Nicholas Hale and Lloyd N. Trefethen. New quadrature formulas from conformal maps. SIAM Journal on Numerical Analysis, 46(2):930–948, 2008. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hale:2012:CNQ

[HT12] Nicholas Hale and Lloyd N. Trefethen. Chebfun and numerical quadrature. Science China. Mathematics, 55(9):1749–1760, 2012. ISSN 1674-7283 (print), 1869-1862 (electronic).

Hashemi:2017:CTD

[HT17] Behnam Hashemi and Lloyd N. Trefethen. Chebfun in three dimensions. SIAM Journal on Scientific Computing, 39(5): C341–C363, ???? 2017. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Huybrechs:2023:AIE

[HT23] Daan Huybrechs and Lloyd N. Trefethen. AAA interpolation of equispaced data. BIT Numerical Mathematics, 63(2): 1–19, June 2023. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10543-023-00959-x.

Huybrechs:2024:SFM

[HT24] Daan Huybrechs and Lloyd N.

Trefethen. Sigmoid functions, multiscale resolution of singularities, and hp-mesh refinement. SIAM Review, 66(4): ??, ???? 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://epubs.siam.
org/doi/10.1137/23M1556629.

Hayashi:1990:CT

[HTG90] Eric Hayashi, Lloyd N. Trefethen, and Martin H. Gutknecht. The CF table. Constructive Approximation, 6(2):195–223, 1990. ISSN 0176-4276 (print), 1432-0940 (electronic).

Higham:1998:NAP

[HWG98] D. J. (Desmond J.) Higham, G. A. Watson, and D. F. (David Francis) Griffiths, editors. Numerical analysis 1997: proceedings of the 17th Dundee Biennial Conference, June 24– 27, 1997, volume 380 of Pitman Research Notes in Mathematics. Longman Scientific and Technical, Harlow, Essex, 1998. ISBN 0-582-31261-2 (paperback). ISSN 0269-3674. LCCN QA297 .D85 1997.

Iserles:1999:AN

[Ise99] Arieh Iserles. Acta numerica 1999, volume 8 of Acta Numerica. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-77088-2 (hardcover). ISSN 0962-4929 (print), 1474-0508 (electronic). 295 pp. LCCN QA279 1999.

Jonsson:1998:NAL

[JT98] G. F. Jønsson and L. N. Trefethen. A numerical analyst looks at the “cutoff phenomenon” in card shuffling and other Markov chains. In Higham et al. [HWG98], pages 150– 178. ISBN 0-582-31261-2 (paperback). ISSN 0269-3674. LCCN QA297 .D85 1997.

Javed:2014:TRE

[JT14] Mohsin Javed and Lloyd N. Trefethen. A trapezoidal rule error bound unifying the Euler– Maclaurin formula and geometric convergence for periodic functions. Proceedings of the Royal Society A: Mathematical, Physical, & Engineering Sciences, 470(2161):20130571, 2014. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Javed:2016:EMG

[JT16] Mohsin Javed and Lloyd N. Trefethen. Euler–Maclaurin and Gregory interpolants. Num. Math, 132(1):201–216, January 2016. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s00211-015-0713-x.

Keller:2002:LEH

[Kel02] Joseph B. Keller. Letter to the Editor: The hundreddollar, hundred-digit challenge.

SIAM News, 35(10):??, December 2002. ISSN 0036-1437.

URL http://www.siam.org/
siamnews/12-02/letter.htm.

Kern:2002:SHD

[Ker02a] Michel Kern. Solution to the “Hundred-dollar, Hundred-digit Challenge”. World-Wide Web document., May 17, 2002. URL

http://www-rocq.inria.fr/~
kern/Challenge/challenge.

pdf. See [Tre02b, Tre02a, Tre02c].

Kern:2002:SSH

[Ker02b]  Michel Kern.  Solution to

the SIAM “Hundred-dollar, Hundred-digit Challenge”. Report 4472, INRIA, Unité de recherche INRIA Rocquencourt, Domaine de Voluceau, Rocquencourt, BP 105, 78153 Le Chesnay Cedex, France, May 2002. 21 pp. URL

http://www-rocq.inria.fr/
who/Michel.Kern/Challenge/
RR-challenge.pdf;   http:/
/www-rocq.inria.fr/~kern/
Challenge/challenge.html;
 http://www.lacim.uqam.ca/
~plouffe/OEIS/citations/RR-

4472.pdf. See [Tre02b, Tre02a, Tre02c].

Kirchgassner:1996:IPT

[KMM96] Klaus Kirchgassner, Oskar Mahrenholtz, and Reinhard Mennicken, editors. ICIAM 95: proceedings of the Third International Congress on Industrial and Applied Mathematics held in Hamburg, Germany, July 3– 7, 1995, volume 87 of Mathematical research. Akademie Verlag, Berlin, Germany, 1996. ISBN 3-05-501682-3. LCCN QA1 .I73 1995.

Kassam:2005:FOT

[KT05] Aly-Khan Kassam and Lloyd N. Trefethen. Fourth-order timestepping for stiff PDEs. SIAM Journal on Scientific Computing, 26(4):1214–1233, July 2005. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41063.

Kaplan:2002:BSS

[KW02] Danny Kaplan and Stan Wagon. Brief solutions to the SIAM Hundred-Dollar, Hundred-Digit Challenge. World-Wide Web document., June 27, 2002.

                                        URL http://stanwagon.com/
                                        wagon/Misc/HTMLLinks/SIAMchallenge_

2.html. See [Tre02b, Tre02a,

Tre02c].

Laurie:2002:STH

[Lau02] Dirk Laurie. Solutions to Trefethen’s hundred-dollar, hundreddigit challenge problems. World-Wide Web document., May 23,

                                        2002. URL http://dip.sun.
                                        ac.za/~laurie/trefethen-challenge/
                                        ;   http://dip.sun.ac.za/~
                                        laurie/trefethen-challenge/
                                        howsolved.pdf. See [Tre02b,

Tre02a, Tre02c].

LeVeque:1984:RCK

[LT84] Randall J. LeVeque and Lloyd N. Trefethen. On the resolvent condition in the Kreiss Matrix Theorem. BIT, 24(4):584–591, 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

LeVeque:1988:FAS

[LT88] Randall J. LeVeque and Lloyd N.

Trefethen. Fourier analysis of the SOR iteration. IMA J. Numer. Anal., 8(3):273–279, 1988. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

LeVeque:1985:PSS

[LTS85] J. LeVeque, Lloyd N. Trefethen, and James C. Smith. Problems and solutions: Solutions of advanced problems: 6462. Amer. Math. Monthly, 92(10):740–741, December 1985. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DMLT84].

Morton:1988:NMF

[MB88] K. W. Morton and M. J.

(Michael John) Baines, editors. Numerical methods for fluid dynamics III: based on the proceedings of a conference organized by the Institute for Computational Fluid Dynamics of the Universities of Oxford and Reading in association with the Institute of Mathematics and [its] Applications on numerical methods for fluid dynamics, held in Oxford in March 1988, volume 17 of The Institute of Mathematics and Its Applications conference, new series. Clarendon Press, Oxford, UK, 1988. ISBN 0-19-853632-

1. LCCN TA357 .N8726 1988.

US$75.00.

Medley:2002:SC

[MBvG02] Brian Medley, Bernard B. Beard, and Marijke van Gans. The 2002 SIAM challenge. World-Wide Web document.,

                                        2002.  URL http://www.
                                        maxwellian.demon.co.uk/~marijke/
                                        SIAM2002/;   http://www.maxwellian.
                                        demon.co.uk/~marijke/SIAM2002/
                                        siamscim.pdf. See [Tre02b,

Tre02a, Tre02c].

Mason:1987:AAB

[MC87] J. C. Mason and M. G. Cox, editors. Algorithms for approximation: based on the proceedings of the IMA Conference on Algorithms for the Approximation of Functions and Data, held at the Royal Military College of Science, Shrivenham, July 1985, volume 10 of The Institute of Mathematics and Its Applications conference series, new series. Clarendon Press, Oxford, UK, 1987. ISBN 0-19-853612-

7. LCCN QA221 .A5361 1987;

QA221 .I47 1985. US$90.

Mason:1990:AAI

[MC90] J. C. Mason and M. G. Cox, editors. Algorithms for approximation II: based on the proceedings of the Second International Conference on Algorithms for Approximation, held at Royal Military College of Science, Shrivenham, July 1988. Chapman and Hall, London, England, 1990.

ISBN 0-412-34580-3. LCCN QA221 .I54 1988.

Mattics:1985:APS

[MN85] L. E. Mattics and Harald Niederreiter. Advanced problems and solutions: Solutions: 6461. Amer. Math. Monthly, 92 (9):672, November 1985. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DMLT84].

McInturff:2002:SHD

[MS02] Kim McInturff and Peter S. Simon. Solutions for the hundreddollar, hundred-digit challenge. World-Wide Web document., April 29, 2002. URL http:/

/www.vcnet.com/~simonp/;
http://www.vcnet.com/~simonp/
solutions.pdf. See [Tre02b,

Tre02a, Tre02c].

Meseguer:2003:LPF

[MT03] Á. Meseguer and L. N. Trefethen. Linearized pipe flow to Reynolds number 10^{7}. Journal of Computational Physics, 186(1):178–197, March 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0021999103000299.

Nachtigal:1992:HFN

[NRT92a] Noël M. Nachtigal, Satish C.

Reddy, and Lloyd N. Trefethen. How fast are nonsymmetric matrix iterations? SIAM Journal on Matrix Analysis and Applications, 13(3):778–795, July 1992.

CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). Iterative methods in numerical linear algebra (Copper Mountain, CO, 1990).

Nachtigal:1992:HGA

[NRT92b] Noël M. Nachtigal, Lothar Reichel, and Lloyd N. Trefethen. A hybrid GMRES algorithm for nonsymmetric linear systems. SIAM Journal on Matrix Analysis and Applications, 13(3):796– 825, July 1992. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). Iterative methods in numerical linear algebra (Copper Mountain, CO, 1990).

Nakatsukasa:2018:AAR

[NST18] Yuji Nakatsukasa, Olivier Sète, and Lloyd N. Trefethen. The AAA algorithm for rational approximation. SIAM Journal on Scientific Computing, 40(3): A1494–A1522, ???? 2018. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Nakatsukasa:2018:RA

[NT18] Yuji Nakatsukasa and Lloyd N. Trefethen. Rational approximation of x^{n}. Proceedings of the American Mathematical Society, 146(12):5219–5224, 2018. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Nakatsukasa:2020:ARC

[NT20] Yuji Nakatsukasa and Lloyd N. Trefethen. An algorithm for real and complex rational minimax approximation. SIAM Journal on Scientific Computing, 42(5): A3157–A3179, ???? 2020. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Nakatsukasa:2021:RLA

[NT21] Yuji Nakatsukasa and Lloyd N.

Trefethen. Reciprocal-log approximation and planar PDE solvers. SIAM Journal on Numerical Analysis, 59(6):2801– 2822, ???? 2021. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pachon:2010:PSC

[PPT10] Ricardo Pachón, Rodrigo B.

Platte, and Lloyd N. Trefethen. Piecewise-smooth chebfuns. IMA J. Numer. Anal., 30(4):898–916, October 2010. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http://

imajna.oxfordjournals.org/
content/30/4/898.full.pdf+
html.

Pachon:2009:BRA

[PT09] Ricardo Pachón and Lloyd N.

Trefethen. Barycentric–Remez algorithms for best polynomial approximation in the chebfun system. BIT Numerical Mathematics, 49(4):721–741, December 2009. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-

3835&volume=49&issue=4&spage=
721.

Platte:2010:CNK

[PT10] R. B. Platte and L. N. Trefethen. Chebfun: a new kind of numerical computing. In Alistair D. Fitt, John Norbury, Eddie Wilson, and Hilary Ockendon, editors, Progress in industrial mathematics at ECMI 2008, volume 15 of Mathematics in Industry, pages 69– 87. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-12110-1.

Platte:2011:IFS

[PTK11] Rodrigo B. Platte, Lloyd N. Trefethen, and Arno B. J. Kuijlaars. Impossibility of fast stable approximation of analytic functions from equispaced samples. SIAM Review, 53 (2):308–318, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/1/
siread/v53/i2/p308_s1.

Reddy:1990:LSF

[RT90] Satish C. Reddy and Lloyd N. Trefethen. Lax-stability of fully discrete spectral methods via stability regions and pseudoeigenvalues. Comput. Meth. Appl. Mech. Engin., 80(1–3): 147–164, 1990. CODEN CM-MECC. ISSN 0045-7825, 0374-2830. Spectral and high order methods for partial differential equations (Como, 1989).

Reddy:1992:SML

[RT92a] Satish C. Reddy and Lloyd N.

Trefethen. Stability of the method of lines. Num. Math, 62(2):235–267, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reichel:1992:EPE

[RT92b] Lothar Reichel and Lloyd N.

Trefethen. Eigenvalues and pseudo-eigenvalues of Toeplitz matrices. Linear Algebra and its Applications, 162/164:153– 185, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Directions in matrix theory (Auburn, AL, 1990).

Reddy:1994:PCD

[RT94] Satish C. Reddy and Lloyd N.

Trefethen. Pseudospectra of the convection-diffusion operator. SIAM Journal on Applied Mathematics, 54(6):1634– 1649, December 1994. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24698.

Richardson:2011:SFA

[RT11] Mark Richardson and Lloyd N.

Trefethen. A sinc function analogue of Chebfun. SIAM Journal on Scientific Computing, 33 (5):2519–2535, ???? 2011. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (elec-

tronic). URL http://epubs.
siam.org/sisc/resource/1/
sjoce3/v33/i5/p2519_s1.

Reddy:1993:PCD

[RTP93] Satish C. Reddy, Lloyd N. Trefethen, and D. Pathria. Pseudospectrum of the convection– diffusion operator. Technical Report CTC 93 TR126, Advanced Computing Research Institute, Cornell Theory Center, Cornell University, Ithaca, New York, 1993. ???? pp.

Schleicher:2011:IMC

[SL11] Dierk Schleicher and Malte Lackmann, editors. An Invitation to Mathematics: From Competitions to Research. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-19532-6 (soft cover), 3-642-19533-4 (e-book). xiv + 220 pp. LCCN ????

Schleicher:2013:EME

[SL13] Dierk Schleicher and Malte Lackmann. Eine Einladung in die Mathematik: Einblicke in aktuelle Forschung. (German) [An invitation to mathematics: insights into current research]. Springer Spektrum, Heidelberg, Germany, 2013. ISBN 3-642-25797-6, 3-642-25798-4 (e-book). xv + 228 pp. LCCN QA37.3 .E3615 2013eb. URL http:

                                        //public.eblib.com/choice/
                                        publicfullrecord.aspx?p=1082421;
                                         http://site.ebrary.com/id/
                                        10640332.

Saff:1993:FCA

[SS93] E. B. Saff and Arthur David Snider. Fundamentals of complex analysis for mathematics, science, and engineering. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1993. ISBN 0-13-327461-6. xi + 468 pp. LCCN QA300 .S18 1993. US$48.75.

Saff:2003:FCA

[SS03] Edward B. Saff and Arthur David Snider. Fundamentals of Complex Analysis: with Applications to Engineering and Science. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2003. ISBN 0-13-907874-6, 0-13-017968-X (International ed.). xi + 511 pp. LCCN QA300 .S18 2003. With an appendix by Lloyd N. Trefethen and Tobin A. Driscoll.

Schmelzer:2007:CGF

[ST07] Thomas Schmelzer and Lloyd N.

Trefethen. Computing the gamma function using contour integrals and rational approximations. SIAM Journal on Numerical Analysis, 45(2):558–571, ???? 2007. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Schmelzer:2007:EMF

[ST08] Thomas Schmelzer and Lloyd N.

Trefethen. Evaluating matrix functions for exponential integrators via Carathéodory–Fejér approximation and contour integrals. Electronic Transactions on Numerical Analysis, 29:1– 18, 2007/2008. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

                                        /etna.mcs.kent.edu/vol.29.
                                        2007-2008/pp1-18.dir/pp1-
                                        18.pdf.

Suli:2010:BMB

[STW10] Endre Süli, Nick Trefethen, and Andy Wathen. Bill Morton’s 80th birthday conference. Mathematics Today, 46(4):174–175, 2010. ISSN 1361-2042.

Specht:2002:HSH

[SWRG02] Eckard Specht, Martin Wohlgemuth, Clemens Raab, and Lutz Gehlen. How to solve the “A Hundred-Dollar, Hundred-Digit Challenge”. World-Wide Web document., 2002.

                                        URL http://matheplanet.
                                        com/matheplanet/nuke/html/
                                        matroid/siam/siam_results.

pdf. See [Tre02b, Tre02a, Tre02c].

Trefethen:1984:ADP

[TA84] Lloyd N. Trefethen and Z. Angew. Analysis and design of polygonal resistors by conformal mapping. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 35(5):692–704, 1984. CODEN AZMPA8. ISSN 0044-2275 (print), 1420-9039 (electronic).

Trefethen:1997:LDM

[TB97a] L. N. Trefethen and J. S. Baggett. Low-dimensional models of subcritical transition to turbulence. Physics of Fluids, 9 (4):1043–1053, ???? 1997. CO-DEN PHFLE6. ISSN 1070-6631.

Trefethen:1997:NLA

[TB97b] Lloyd N. Trefethen and David Bau, III, editors. Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-361-7. xii + 361 pp. LCCN QA184 .T74 1997.

Trefethen:2006:CEP

[TB06] Lloyd N. Trefethen and Timo Betcke. Computed eigenmodes of planar regions. In Recent advances in differential equations and mathematical physics, volume 412 of Contemp. Math., pages 297–314. Amer. Math. Soc., Providence, RI, USA, 2006.

Trefethen:2022:NLA

[TB22] Lloyd N. Trefethen and David Bau, III. Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, anniversary edition, 2022. ISBN 1-61197-715-0; 1-61197-716-9. xvi + 370 pp. With a foreword by James G. Nagy.

Trefethen:2018:EO

[TBD18] Lloyd N. Trefethen, Ásgeir Birkisson, and Tobin A. Driscoll. Exploring ODEs, volume 157. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2018. ISBN 1-61197-515-8 (hardcover). vii + 335 pp. LCCN QA371 .T67 2018.

Trefethen:2004:WPP

[TC04] Lloyd N. Trefethen and S. J. Chapman. Wave packet pseudomodes of twisted Toeplitz matrices. Comm. Pure Appl. Math., 57(9):1233–1264, 2004. CO-DEN CPAMAT. ISSN 0010-3640 (print), 1097-0312 (electronic).

Trefethen:2001:SPL

[TCE01] Lloyd N. Trefethen, Marco Contedini, and Mark Embree. Spectra, pseudospectra, and localization for random bidiagonal matrices. Comm. Pure Appl. Math., 54(5):595–623, 2001. CODEN CPAMAT. ISSN 0010-3640 (print), 1097-0312 (electronic).

Trefethen:1998:SCM

[TD98] Lloyd N. Trefethen and Tobin A. Driscoll. Schwarz– Christoffel mapping in the computer era. In Fischer et al. [FSR98], pages 533–542. ISSN 1431-0635. LCCN QA1 .I61 1998. Extra volume III.

Trefethen:19xx:PTM

[TDTxx] L. N. Trefethen, T. A. Driscoll, and K.-C. Toh. From potential theory to matrix iterations in six steps. SIAM Journal on Matrix Analysis and Applications, ?? (??):??, ???? 19xx. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). To appear.

Trefethen:2005:SPB

[TE05] Lloyd N. Trefethen and Mark Embree. Spectra and Pseudospectra: the Behavior of Nonnormal Matrices and Operators. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11946-5. xviii + 606 pp. LCCN QA320 .T67 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0654/2005046573-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0654/
2005046573-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0654/2005046573-t.html.

The behavior of nonnormal matrices and operators.

Trefethen:1983:NBR

[TG83a] Lloyd N. Trefethen and M. H.

Gutknecht. Nonuniqueness of best rational Chebyshev approximations on the unit disk. Journal of Approximation Theory, 39(3):275–288, 1983. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Trefethen:1983:RCCa

[TG83b] Lloyd N. Trefethen and M. H.

Gutknecht. Real and complex Chebyshev approximation on the unit disk and interval. Bull. Amer. Math. Soc., 8(3):455–458, 1983. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Trefethen:1983:CFM

[TG83c] Lloyd N. Trefethen and Martin H. Gutknecht. The Carathéodory–Fejér method for real rational approximation. SIAM Journal on Numerical Analysis, 20(2):420–436, April 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Trefethen:1983:RCCb

[TG83d] Lloyd N. Trefethen and Martin H. Gutknecht. Real and complex Chebyshev approximation on the unit disk and interval. In Chui et al. [CSW83], pages 712– 722. ISBN 0-12-174580-5. LCCN QA297.5 .I54 1983.

Trefethen:1983:RVC

[TG83e] Lloyd N. Trefethen and Martin H. Gutknecht. Real vs. complex rational Chebyshev approximation on an interval. Transactions of the American Mathematical Society, 280(2):555–561, 1983. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Trefethen:1984:RVC

[TG84] Lloyd N. Trefethen and M. H. Gutknecht. Real vs. complex rational Chebyshev approximation on complex domains. In Collatz et al. [CMW82], pages 87– 97. ISBN 3-7643-1304-8. LCCN QA221 .T129n 1981; QA297.5 .N85.

Trefethen:1985:CDR

[TG85] Lloyd N. Trefethen and Martin H. Gutknecht. On convergence and degeneracy in rational Padé and Chebyshev approximation. SIAM Journal on Mathematical Analysis, 16(1): 198–210, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trefethen:1987:PSP

[TG87] Lloyd N. Trefethen and M. H. H.

Gutknecht. Padé, stable Padé, and Chebyshev–Padé approximation. In Mason and Cox [MC87], pages 227–264. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Trefethen:1997:DPM

[TG97] Lloyd N. Trefethen and A. Greenbaum. Do the pseudospectra of a matrix determine its behavior? Linear Algebra and its Applications, ??:??, 1997. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). To appear. The first author reports that the paper has been accepted, subject to minor revisions.

Trefethen:1986:WPA

[TH86a] Lloyd N. Trefethen and L. Halpern.

Well-posedness of absorbing boundary conditions and oneway wave equations. Mathematics of Computation, 47(176): 421–435, 1986. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trefethen:1986:WPO

[TH86b] Lloyd N. Trefethen and Laurence Halpern. Well-posedness of one-way wave equations and absorbing boundary conditions.

Mathematics of Computation, 47(176):421–435, 1986. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trefethen:1988:IPA

[TH88] Lloyd N. Trefethen and L. H. Howell. Ill-posedness of absorbing boundary conditions for migration. Geophysics, 53:593– 603, 1988. CODEN GPYSA7. ISSN 0016-8033 (print), 1942-2156 (electronic).

Trefethen:1997:EMI

[TH97] L. N. Trefethen and V. E. Howle. Eigenvalues and musical instruments. Technical Report ????, ????, ????, August 1997.

Trefethen:1990:CT

[THG90] Lloyd N. Trefethen, E. Hayashi, and M. H. Gutknecht. The CF table. Constructive Approximation, 6(2):195–223, 1990. ISSN 0176-4276 (print), 1432-0940 (electronic).

Trefethen:1998:NAL

[TJ98] L. N. Trefethen and G. F. Jønsson. A numerical analyst looks at the ‘cutoff phenomenon’ in card shuffling and other Markov chains. In Higham et al. [HWG98], page ?? ISBN 0-582-31261-2 (paperback). ISSN 0269-3674. LCCN QA297 .D85 1997.

Trefethen:1984:PSA

[TLMD84] Lloyd N. Trefethen, Randall J. LeVeque, L. E. Mattics, and L. Richard Duffy. Problems and solutions: Advanced problems: 6460–6462. Amer. Math. Monthly, 91(6):371, 1984. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Trefethen:1999:MMM

[TMC^{+}99] Anne E. Trefethen, Vijay S.

Menon, Chi-Chao Chang, Grzegorz J. Czajkowski, Chris Myers, and Lloyd N. Trefethen. MultiMATLAB: MATLAB on multiple processors. Technical report, Cornell Theory Center and Computer Science Department, Cornell University, Ithaca, NY, August 19, 1999. URL http:

//www.cs.cornell.edu/Info/
People/lnt/multimatlab.html.

Trefethen:2021:ENC

[TNW21] Lloyd N. Trefethen, Yuji Nakatsukasa, and J. A. C. Weideman. Exponential node clustering at singularities for rational approximation, quadrature, and PDEs. Num. Math, 147(1):227–254, January 2021. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.com/article/
10.1007/s00211-020-01168-
2;   https://link.springer.
com/content/pdf/10.1007/s00211-
020-01168-2.pdf.

Trefethen:1990:SUQ

[TP90] Lloyd M. Trefethen and Ronald L.

Panton. Some unanswered questions in fluid mechanics. Applied Mechanics Reviews, 43(8): 153–170, 1990. CODEN AM-READ. ISSN 0003-6900 (print), 1088-8535 (electronic).

Trefethen:1992:SML

[TR92a] Lloyd N. Trefethen and S. C. Reddy. Stability of the method of lines. Num. Math, 62(2):235– 267, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Trefethen:1992:EPE

[TR92b] Lloyd N. Trefethen and L. Reichel. Eigenvalues and pseudoeigenvalues of Toeplitz matrices. Linear Algebra and its Applications, 162–164:153–185, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trefethen:1977:CAP

[Tre77] Lloyd N. Trefethen. Chebyshev approximation by polynomials in the complex plane. Undergraduate thesis, Harvard University, Cambridge, MA, USA, May 1977.

Trefethen:1980:CAS

[Tre80a] Lloyd N. Trefethen. Computation and application of Schwarz– Christoffel transformations. In Anonymous [Ano80], pages 165– 171. ISBN ???? LCCN ???? Six microfiches.

Trefethen:1980:ENC

[Tre80b] Lloyd N. Trefethen. Erratum: “Numerical computation of the Schwarz–Christoffel transformation”. SIAM Journal on Scientific and Statistical Computing, 1(2):302, 1980. CODEN SI-JCD4. ISSN 0196-5204. See [Tre80c].

Trefethen:1980:NCS

[Tre80c]  Lloyd N. Trefethen.  Nu-

merical computation of the Schwarz–Christoffel transformation. SIAM Journal on Scientific and Statistical Computing, 1(1):82–102, 1980. CODEN SI-JCD4. ISSN 0196-5204. See erratum [Tre80b].

Trefethen:1981:CAS

[Tre81a] Lloyd N. Trefethen. Computer application of the Schwarz– Christoffel transformation. In Butzer and Fehér [BF81], pages 263–274. ISBN 3-7643-1162-2. LCCN QA7 I57 1981; QA 7 .I57 1981; QA76 .I556

1981. URL https://link.
springer.com/book/10.1007/
978-3-0348-5452-8.

Trefethen:1981:NCC

[Tre81b] Lloyd N. Trefethen. Nearcircularity of the curve in complex Chebyshev approximation. Journal of Approximation Theory, 31(4):344–367, 1981. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Trefethen:1981:NCE

[Tre81c] Lloyd N. Trefethen. Nearcircularity of the error curve in complex Chebyshev approximation. Journal of Approximation Theory, 31(4):344–367, 1981. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Trefethen:1981:RCA

[Tre81d] Lloyd N. Trefethen. Rational Chebyshev approximation on the unit disk. Num. Math, 37(2):297–320, 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Trefethen:1982:GVF

[Tre82a] Lloyd N. Trefethen. Group velocity in finite difference schemes. SIAM Review, 24 (2):113–136, April 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1982:WPS

[Tre82b] Lloyd Nicholas Trefethen. Wave Propagation and Stability for Finite Difference Schemes. Ph.D. thesis, Stanford University, Stanford, CA, USA, 1982. 208 pp.

Trefethen:1983:CAU

[Tre83a] Lloyd N. Trefethen. Chebyshev approximation on the unit disk. In Werner et al. [W^{+}83], pages 309–323. ISBN 90-277-1571-8. LCCN QA300 .N29 1982.

Trefethen:1983:CEC

[Tre83b] Lloyd N. Trefethen. Circularity of the error curve and sharpness of the CF method in complex Chebyshev approximation. SIAM Journal on Numerical Analysis, 20(6):1258–1263, December 1983. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Trefethen:1983:GVI

[Tre83c] Lloyd N. Trefethen. Group velocity interpretation of the stability theory of Gustafsson, Kreiss, and Sundström. Journal of Computational Physics, 49(2):199–217, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0021999183901237.

Trefethen:1984:AAR

[Tre84a] Lloyd N. Trefethen. The asymptotic accuracy of rational best approximations to e^{z} on a disk. Journal of Approximation Theory, 40(4):380–383, 1984. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Trefethen:1984:BRF

[Tre84b] Lloyd N. Trefethen. Book review: Fourier Analysis of Numerical Approximations of Hyperbolic Equations (R. Vichnevetsky and J. B. Bowles). SIAM Review, 26(3):439–441, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1984:IDM

[Tre84c] Lloyd N. Trefethen. Instability of difference models for hyperbolic initial-boundary value problems. Comm. Pure Appl. Math., 37(3):329–367, 1984.

CODEN CPAMAT. ISSN 0010-3640 (print), 1097-0312 (electronic).

Trefethen:1984:IFD

[Tre84d] Lloyd N. Trefethen. Instability of finite difference models for hyperbolic initial boundary value problems. Comm. Pure Appl. Math., 24(3):329– 367, 1984. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Trefethen:1984:IDD

[Tre84e] Lloyd N. Trefethen. On ℓ^{p}- instability and dispersion at discontinuities in finite difference schemes. In Vichnevetsky and Stepleman [VS84b], page ?? ISBN ???? LCCN ????

Trefethen:1984:SBE

[Tre84f] Lloyd N. Trefethen. Square blocks and equioscillation in the Padé, Walsh, and CF tables. Lecture Notes in Mathematics, 1105:170–181, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/
10.1007/BFb0072410/.

Trefethen:1985:SFD

[Tre85a] Lloyd N. Trefethen. Stability of finite-difference models containing two boundaries or interfaces. Mathematics of Computation, 45(172):279–300, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Trefethen:1985:SHF

[Tre85b] Lloyd N. Trefethen. Stability of hyperbolic finite-difference models with one or two boundaries. In Engquist et al. [EOS85], pages 311–326. ISBN 0-8218-1122-3. LCCN QA901 .L37 1985.

Trefethen:1985:TMG

[Tre85c] Lloyd N. Trefethen. Three mysteries of Gaussian elimination. SIGNUM Newsletter, 20 (4):2–5, October 1985. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Trefethen:1986:DDS

[Tre86a] Lloyd N. Trefethen. Dispersion, dissipation, and stability. In Griffiths and Watson [GW86], pages 200–219. ISBN 0-470-20669-1, 0-582-98897-7. LCCN QA297 .N8251 1986. The 11th Dundee Biennial Conference on Numerical Analysis, held at the University of Dundee on 25–28 June 1985.

Trefethen:1986:MPC

[Tre86b] Lloyd N. Trefethen. MATLAB programs for CF approximation. In Chui et al. [CSW86], page ?? ISBN 0-12-174581-3. LCCN QA221 .S92 1986; QA221.S92. Proceedings of the Fifth International Symposium on Approximation Theory, held at Texas A and M University on January 13–17, 1986.

Trefethen:1986:NCM

[Tre86c] Lloyd N. Trefethen, editor. Numerical Conformal Mapping. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-87855-6. iv + 269 pp. LCCN QA1.J54. Reprint of J. Comput. Appl. Math. 14 (1986), no. 1–2.

Trefethen:1986:P

[Tre86d]  Lloyd N. Trefethen.  Pref-

ace. Journal of Computational and Applied Mathematics, 14 (1–2):1–6, February 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0377042786901275.

Trefethen:1986:SIN

[Tre86e] Lloyd N. Trefethen, editor. Special issue on numerical conformal mapping, volume 14(1–2) of Journal of Computational and Applied Mathematics. Elsevier Science Publishers, Amsterdam, The Netherlands, 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). viii + 269 pp.

Trefethen:1988:LSV

[Tre88]  Lloyd N. Trefethen.  Lax-

stability vs. eigenvalue stability of spectral methods. In Morton and Baines [MB88], pages 237– 253. ISBN 0-19-853632-1. LCCN TA357 .N8726 1988. US$75.00.

Trefethen:1989:SCM

[Tre89a] Lloyd N. Trefethen. Schwarz– Christoffel mapping in the 1980’s. Numerical Analysis Report 89-1, Dept. of Mathematics, MIT, Cambridge, MA, USA, 1989.

Trefethen:1989:SUG

[Tre89b] Lloyd N. Trefethen. SCPACK user’s guide. Numerical Analysis Report 89-2, Dept. of Mathematics, MIT, Cambridge, MA, USA, 1989. (An earlier edition appeared as an ICASE internal report in 1983.).

Trefethen:1990:ATN

[Tre90] Lloyd N. Trefethen. Approximation theory and numerical linear algebra. In Mason and Cox [MC90], pages 336–360. ISBN 0-412-34580-3. LCCN QA221 .I54 1988.

Trefethen:1992:DNA

[Tre92a]  Lloyd N. Trefethen.  The

definition of numerical analysis. SIAM News, 25, November 6, 1992. ISSN 0036-

1437. URL http://webs.um.
es/eliseo/um/uploads/Main/
TrefethendefNA.pdf. Reprinted

in the Bulletin of the Institute for Mathematics and its Applications, 1993.

Trefethen:1992:PM

[Tre92b] Lloyd N. Trefethen. Pseudospectra of matrices. In Griffiths and Watson [GW91], pages 234– 266. ISBN 0-582-08908-5. LCCN QA297 .D85 1991.

Trefethen:1993:NCC

[Tre93] Lloyd N. Trefethen. Numerical construction of conformal maps. In Fundamentals of complex analysis for mathematics, science, and engineering [SS93], page ?? ISBN 0-13-327461-6. LCCN QA300 .S18 1993. US$48.75.

Trefethen:1994:SPP

[Tre94] L. N. Trefethen. Spectra and pseudospectra for pipe Poiseuille flow. IPS Research Report 94-16, Interdisciplinary Project Center for Supercomputing, ETH-Zurich, Zurich, Switzerland, ???? 1994.

Trefethen:1996:PLO

[Tre96] L. N. Trefethen. Pseudospectra of linear operators. In Kirchgassner et al. [KMM96], pages 401–434. ISBN 3-05-501682-3. LCCN QA1 .I73 1995. Also in [Tre97].

Trefethen:1997:PLO

[Tre97] Lloyd N. Trefethen. Pseudospectra of linear operators. SIAM Review, 39(3):383–406, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29528.

Trefethen:1999:CP

[Tre99a] Lloyd N. Trefethen. Computation of pseudospectra. Acta Numerica, 8:247–295, 1999. CO-DEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Trefethen:1999:SP

[Tre99b] Lloyd N. Trefethen. Spectra and pseudospectra. In Mark Ainsworth, Jeremy Levesley, and Marco Marletta, editors, The Graduate Student’s Guide to Numerical Analysis ’98, pages 217–250. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999.

Trefethen:1999:SPB

[Tre99c] Lloyd N. Trefethen. Spectra and pseudospectra. The behaviour of non-normal matrices and operators. In Ainsworth et al. [ALM99], pages 217–250. ISBN 3-540-65752-5. ISSN 0179-3632. LCCN QA297 .E57 1998.

Trefethen:1999:SRa

[Tre99d] Nick Trefethen. Survey and review. SIAM Review, 41 (1):1–2, March 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97019.

Trefethen:1999:SRb

[Tre99e] Nick Trefethen. Survey and review. SIAM Review, 41 (2):197, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97024.

Trefethen:1999:SRc

[Tre99f] Nick Trefethen. Survey and review. SIAM Review, 41(3): 415, September 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97029.

Trefethen:1999:SRd

[Tre99g] Nick Trefethen. Survey and review. SIAM Review, 41(4): 635, December 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97035.

Trefethen:2000:SMM

[Tre00a] Lloyd N. Trefethen. Spectral Methods in MATLAB. Software, environments, tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-465-6 (paperback), 0-89871-959-3 (ebook). xvi + 165 pp. LCCN QA377 .T65 2000.

Trefethen:2000:SRa

[Tre00b] Nick Trefethen. Survey and review. SIAM Review, 42 (1):1, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97041.

Trefethen:2000:SRb

[Tre00c] Nick Trefethen. Survey and review. SIAM Review, 42 (2):159, June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97050.

Trefethen:2000:SRIa

[Tre00d] Nick Trefethen. Survey and review — introduction. SIAM Review, 42(3):367, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97051.

Trefethen:2000:SRIb

[Tre00e] Nick Trefethen. Survey and review — introduction. SIAM Review, 42(4):553, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97056.

Trefethen:2001:SRa

[Tre01a] Nick Trefethen. Survey and review. SIAM Review, 43 (1):1, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97061.

Trefethen:2001:SRb

[Tre01b] Nick Trefethen. Survey and review. SIAM Review, 43 (2):233, June 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97066.

Trefethen:2001:SRc

[Tre01c] Nick Trefethen. Survey and review. SIAM Review, 43(3): 411, September 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97071.

Trefethen:2001:SRd

[Tre01d] Nick Trefethen. Survey and review. SIAM Review, 43(4): 583, December 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97076.

Trefethen:2002:DCC

[Tre02a] Lloyd N. Trefethen. The $100, 100-digit challenge: Chastened challenge sponsor: “I misjudged”. SIAM News, 35(6): 4, July/August 2002. ISSN 0036-1437. URL http://

                                        www.siam.org/siamnews/07-
                                        02/challengeupdate.pdf.

Trefethen:2002:HDH

[Tre02b] Lloyd N. Trefethen. A hundreddollar, hundred-digit challenge. SIAM News, 35(1):??, January 2002. ISSN 0036-1437.

                                        URL http://dip.sun.ac.za/~
                                        laurie/trefethen-challenge/
                                        challenge.ps;   http://mathworld.
                                        wolfram.com/Hundred-DollarHundred-
                                        DigitChallengeProblems.html;
                                         http://spiff.rit.edu/classes/
                                        phys317/lectures/intro/hundred_
                                        dollar.pdf;   http://web.
                                        comlab.ox.ac.uk/oucl/work/

nick.trefethen/hundred.html;
http://www.siam.org/siamnews/
01-02/challenge.pdf. See so-

lutions [BLWW04, DHN^{+}02, GI02, KW02, Ker02a, Ker02b, Lau02, MS02, MBvG02, SWRG02, Tre02a, Tre02c]. From the solutions Web site: “Ninety-four teams from twenty-five countries entered the competition. Of these, twenty teams scored 100 points and thus are counted as FIRST PRIZE WINNERS”.

Trefethen:2002:SDD

[Tre02c]  Lloyd N. Trefethen.  The

SIAM 100-dollar, 100-digit challenge. SIAM News, 35(6): 2, July/August 2002. ISSN 0036-1437. URL http://

www.siam.org/siamnews/06-
02/challengedigits.pdf.

Trefethen:2002:SRI

[Tre02d] N. Trefethen. Survey and review — introduction. SIAM Review, 44(4):523, ???? 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2002:SRa

[Tre02e] Nick Trefethen. Survey and review. SIAM Review, 44 (1):1, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97081.

Trefethen:2002:SRb

[Tre02f] Nick Trefethen. Survey and review. SIAM Review, 44 (2):167, June 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97086.

Trefethen:2002:SRc

[Tre02g] Nick Trefethen. Survey and review. SIAM Review, 44(3): 309, September 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97091.

Trefethen:2002:SRd

[Tre02h] Nick Trefethen. Survey and review. SIAM Review, 44(4): 523, December 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97096.

Trefethen:2005:WPP

[Tre05] Lloyd N. Trefethen. Wave packet pseudomodes of variable coefficient differential operators. Proceedings of the Royal Society A: Mathematical, Physical, & Engineering Sciences, 461 (2062):3099–3122, July 13, 2005. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Trefethen:2006:NSP

[Tre06]  Nike Trefethen.  Nonhermi-

tian systems and pseudospectra. In Seminaire: Équations aux Dérivées Partielles. 2005–2006, Sémin. Équ. Dériv. Partielles, pages x + 13. École Polytechnique, Palaiseau, France, 2006.

URL http://www.numdam.org/
item?id=SEDP_2005-2006____
A10_0.

Trefethen:2007:CNF

[Tre07] Lloyd N. Trefethen. Computing numerically with functions instead of numbers. Mathematics in Computer Science, 1 (1):9–19, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=1661-
8270&volume=1&issue=1&spage=
9.

Trefethen:2008:GQB

[Tre08]  Lloyd N. Trefethen.  Is

Gauss quadrature better than Clenshaw–Curtis? SIAM Review, 50(1):67–87, 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/?SIR/50/67/1.

Trefethen:2010:HTQ

[Tre10] Lloyd N. Trefethen. Householder triangularization of a quasimatrix. IMA J. Numer. Anal., 30(4):887–897, October 2010. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http://

imajna.oxfordjournals.org/
content/30/4/887.full.pdf+
html.

Trefethen:2011:SMP

[Tre11a] Lloyd N. Trefethen. Six myths of polynomial interpolation and quadrature. Mathematics Today, 47(4):184–188, 2011. CO-DEN ???? ISSN 1361-2042.

Trefethen:2011:TDP

[Tre11b] Lloyd N. Trefethen. Ten digit problems. In Schleicher and Lackmann [SL11], pages 119– 136. ISBN 3-642-19532-6 (soft cover), 3-642-19533-4 (e-book). LCCN ???? URL http://

link.springer.com/chapter/
10.1007/978-3-642-19533-4_
9.

Trefethen:2011:TIC

[Tre11c] Lloyd N. Trefethen. Trefethen’s index cards: forty years of notes about people, words and mathematics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4360-69-4 (paperback), 981-4458-41-4 (e-book). xv + 368 pp. LCCN ???? URL http:

//www.worldscientific.com/
worldscibooks/10.1142/8207.

Trefethen:2013:ATA

[Tre13a] Lloyd N. Trefethen. Approximation Theory and Approximation Practice. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2013. ISBN 1-61197-239-6 (paperback). viii + 305 pp. LCCN QA221 .T73 2013.

Trefethen:2013:BRV

[Tre13b]  Lloyd N. Trefethen.  Book

review: Visual complex functions: an introduction with phase portraits, by Elias Wegert.

Birkhäuser/Springer Basel AG, Basel, 2012. $59.95. xiv + 360 pp., softcover. ISBN 978-3-0348-0179-9. SIAM Review, 55 (4):791–797, 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2013:ZPG

[Tre13c] Lloyd N. Trefethen. Zehnstellige Probleme. (German) [Ten-digit problems]. In Eine Einladung in die Mathematik: Einblicke in aktuelle Forschung. (German) [An invitation to mathematics: insights into current research] [SL13], pages 121–140. ISBN 3-642-25797-6, 3-642-25798-4 (e-book). LCCN QA37.3 .E3615 2013eb. URL http:

//public.eblib.com/choice/
publicfullrecord.aspx?p=1082421;
 http://site.ebrary.com/id/
10640332.

Trefethen:2015:CNF

[Tre15a] Lloyd N. Trefethen. Computing numerically with functions instead of numbers. Communications of the ACM, 58(10): 91–97, October 2015. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.
acm.org/magazines/2015/10/
192390/fulltext.

Trefethen:2015:PCA

[Tre15b]  Lloyd N. Trefethen.  The

Princeton companion to applied mathematics [book review of MR3380576]. SIAM Review, 57(3):469–473, 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2016:IY

[Tre16a] Lloyd N. Trefethen. Inverse Yogiisms. Notices of the American Mathematical Society, 63 (11):1281–1285, December 2016. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trefethen:2016:SFC

[Tre16b]  Lloyd N. Trefethen.  Sur-

prises of the Faraday cage. SIAM News, 49(6):??, July/ August 2016. ISSN 0036-1437.

                                        URL https://sinews.siam.
                                        org/DetailsPage/TabId/900/
                                        ArtMID/2243/ArticleID/757/
                                        Surprises-of-the-Faraday-
                                        Cage.aspx.

Trefethen:2017:CAI

[Tre17a] Lloyd N. Trefethen. Cubature, approximation, and isotropy in the hypercube. SIAM Review, 59(3):469–491, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2017:MPA

[Tre17b] Lloyd N. Trefethen. Multivariate polynomial approximation in the hypercube. Proceedings of the American Mathematical Society, 145(11):4837–4844, 2017. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Trefethen:2018:SSL

[Tre18] Lloyd N. Trefethen. Series solution of Laplace problems. The ANZIAM Journal, 60(1):1– 26, 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

Trefethen:2020:ATA

[Tre20a] Lloyd N. Trefethen. Approximation theory and approximation practice. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, extended edition, 2020. ISBN 1-61197-593-X. xi + 363 pp.

Trefethen:2020:EPE

[Tre20b] Lloyd N. Trefethen. Eight perspectives on the exponentially ill-conditioned equation εy′′ − xy′ + y = 0. SIAM Review, 62 (2):439–462, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2020:NCM

[Tre20c] Lloyd N. Trefethen. Numerical conformal mapping with rational functions. Computational Methods and Function Theory, 20(3-4):369–387, 2020. ISSN 1617-9447 (print), 2195-3724 (electronic).

Trefethen:2020:QIC

[Tre20d] Lloyd N. Trefethen. Quantifying the ill-conditioning of analytic continuation. BIT Numerical Mathematics, 60(4): 901–915, December 2020. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL https://

link.springer.com/article/
10.1007/s10543-020-00802-7.

Trefethen:2020:RFB

[Tre20e] Lloyd N. Trefethen. Rational functions and beyond. SIAM News, 53(10):1, 3, 2020. ISSN 1557-9573.

Trefethen:2022:AMA

[Tre22a] Lloyd N. Trefethen. An Applied Mathematician’s Apology. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2022. ISBN 1-61197-718-5. ix + 79 pp.

Trefethen:2022:EQF

[Tre22b]  Lloyd N. Trefethen.  Ex-

actness of quadrature formulas. SIAM Review, 64(1): 132–150, ???? 2022. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://epubs.siam.
org/doi/10.1137/20M1389522.

Trefethen:2022:SRT

[Tre22c] Lloyd N. Trefethen. Some recollections of Trefethen and Bau on the occasion of its 25th anniversary. SIAM News, 55(5):6, 2022. ISSN 1557-9573.

Trefethen:2023:NAC

[Tre23a] Lloyd N. Trefethen. Numerical analytic continuation. Japan Journal of Industrial and Applied Mathematics, 40(3):1587– 1636, 2023. CODEN JAPJI7. ISSN 0916-7005 (print), 1868-937X (electronic).

Trefethen:2023:SLE

[Tre23b] Lloyd N. Trefethen. Spectacularly large expansion coefficients in Müntz’s theorem. La Matematica, 2(1):31–36, 2023. ISSN 2730-9657.

Trefethen:2023:BML

[Tre23c] Nick Trefethen. A bifurcation in Moore’s law? SIAM News, 56 (7):2, 2023. ISSN 0036-1437.

Trefethen:1987:ACS

[TS87] L. N. Trefethen and R. S.

Schreiber. Average case stability of Gaussian elimination. Technical Report 88-3, Numerical Analysis, Department of Mathematics, MIT, Cambridge, MA, USA, 1987.

Trefethen:1990:ACS

[TS90] Lloyd N. Trefethen and Robert S.

Schreiber. Average-case stability of Gaussian elimination. SIAM Journal on Matrix Analysis and Applications, 11(3):335– 360, July 1990. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Trefethen:1985:PSS

[TSL85] Lloyd N. Trefethen, James C.

Smith, and J. LeVeque. Problems and solutions: Solutions of advanced problems: 6462. Amer. Math. Monthly, 92(10): 740–741, 1985. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Trefethen:1987:IPS

[TT87] Lloyd N. Trefethen and Manfred R. Trummer. An instability phenomenon in spectral methods. SIAM Journal on Numerical Analysis, 24(5):1008–1023, October 1987. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Toh:1994:PPP

[TT94] Kim-Chuan Toh and Lloyd N. Trefethen. Pseudozeros of polynomials and pseudospectra of companion matrices. Num. Math, 68(3):403–425, 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Toh:1996:CPA

[TT96a] Kim-Chuan Toh and Lloyd N. Trefethen. Calculation of pseudospectra by the Arnoldi iteration. SIAM Journal on Scientific Computing, 17(1):1– 15, January 1996. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). Special issue on iterative methods in numerical linear algebra (Breckenridge, CO, 1994).

Trefethen:1996:MMM

[TT^{+}96b] L. N. Trefethen, A. E. Trefethen, et al. MultiMAT-LAB: MATLAB on multiple processors. Technical Report CTC96TR293, Cornell Theory Center, Cornell University, Ithaca, NY, USA, ???? 1996.

Toh:1998:CPM

[TT98] Kim-Chuan Toh and Lloyd N.

Trefethen. The Chebyshev polynomials of a matrix. SIAM Journal on Matrix Analysis and Applications, 20(2):400– 419, April 1998. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/30373.

Toh:1999:KMT

[TT99] Kim-Chuan Toh and Lloyd N.

Trefethen. The Kreiss matrix theorem on a general complex domain. SIAM Journal on Matrix Analysis and Applications, 21(1):145–165, 1999. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Toh:2000:KMT

[TT00a] Kim-Chuan Toh and Lloyd N.

Trefethen. The Kreiss matrix theorem on a general complex domain. SIAM Journal on Matrix Analysis and Applications, 21(1):145–165, January 2000. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32402.

Trefethen:2000:HMS

[TT00b] L. N. Trefethen and L. M. Trefethen. How many shuffles to randomize a deck of cards? Proceedings of the Royal Society A: Mathematical, Physical, & Engineering Sciences, 456(2002):

2561–2568, October 8, 2000. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Tee:2006:RSC

[TT06] T. W. Tee and Lloyd N. Trefethen. A rational spectral collocation method with adaptively transformed Chebyshev grid points. SIAM Journal on Scientific Computing, 28(5): 1798–1811 (electronic), 2006. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Townsend:2013:ECT

[TT13] Alex Townsend and Lloyd N. Trefethen. An extension of Chebfun to two dimensions. SIAM Journal on Scientific Computing, 35(6):C495–C518, ???? 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Townsend:2015:CAM

[TT15] A. Townsend and L. N. Trefethen. Continuous analogues of matrix factorizations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 471(2173): 20140585, January 8, 2015. CO-DEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL http://rspa.
royalsocietypublishing.org/
content/471/2173/20140585.

Trefethen:1992:PLN

[TTR92] Lloyd N. Trefethen, A. E. Trefethen, and S. C. Reddy. Pseudospectra of the linear Navier– Stokes evolution operator and instability of plane Poiseuille and Couette flows. Technical Report TR 92-1291, Dept. of Comp. Sci., Cornell University, Ithaca, NY, USA, 1992.

Trefethen:1992:NDH

[TTRD92] L. N. Trefethen, A. E. Trefethen, S. C. Reddy, and T. A. Driscoll. A new direction in hydrodynamic stability: beyond eigenvalues. Technical Report CTC92TR115, Cornell Theory Center, Cornell University, Ithaca, NY, USA, ???? 1992.

Trefethen:1993:HSE

[TTRD93] Lloyd N. Trefethen, Anne E.

Trefethen, Satish C. Reddy, and Tobin A. Driscoll. Hydrodynamic stability without eigenvalues. Science, 261(5121): 578–584, July 30, 1993. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Trefethen:1999:SPP

[TTS99] Anne E. Trefethen, Lloyd N.

Trefethen, and Peter J. Schmid. Spectra and pseudospectra for pipe Poiseuille flow. Comput. Meth. Appl. Mech. Engin., 175 (3–4):413–420, 1999. CODEN CMMECC. ISSN 0045-7825, 0374-2830.

Trefethen:1996:MBU

[TV96] L. N. Trefethen and D. Viswanath. Matrix behaviour, unitary reducibility, and Hadamard products. Technical Report TR96-1596, Department of Computer Science, Cornell University, Ithaca, NY, USA, July 1996.

Trefethen:1986:CMS

[TW86] Lloyd N. Trefethen and Ruth J. Williams. Conformal mapping solution of Laplace’s equation on a polygon with oblique derivative boundary conditions. Journal of Computational and Applied Mathematics, 14(1–2): 227–249, February 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/037704278690141X.

Trefethen:1991:TRP

[TW91] Lloyd N. Trefethen and J. A. C. Weideman. Two results on polynomial interpolation in equally spaced points. Journal of Approximation Theory, 65(3):247– 260, 1991. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Trefethen:2006:FTR

[TW06] L. N. Trefethen and J. A. C. Weideman. The fast trapezoid rule in scientific computing. Unpublished manuscript., 2006.

Trefethen:2014:ECT

[TW14] Lloyd N. Trefethen and J. A. C.

Weideman. The exponentially convergent trapezoidal rule. SIAM Review, 56(3):385–458, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2006:TQR

[TWS06] L. N. Trefethen, J. A. C. Weideman, and T. Schmelzer. Talbot quadratures and rational approximations. BIT Numerical Mathematics, 46(3):653–670, September 2006. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=46&issue=3&spage=

653. In memory of Germund Dahlquist (1925–2005).

vanderVorst:2002:CMC

[vdVDE^{+}02] Henk van der Vorst, Iain Duff, Howard Elman, Ronal Freund, Tim Kelley, Seymour Parter, Gerhard Starke, Nick Trefethen, Panayot Vassilevski, Homer Walker, and Olof Widlund. 2000 Copper Mountain Conference. SIAM Journal on Scientific Computing, 23(2):vii, March 2002. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97414.

Vichnevetsky:1984:IFD

[VS84a] R. Vichnevetsky and R. Stepleman, editors. Advances in computer methods for partial differential equations-V: proceedings of the fifth IMACS International Symposium on Computer Methods for Partial Differential Equations, held at Lehigh University, Bethlehem, Pennsylvania, USA, June 19–June 21, 1984. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1984. LCCN QA377.I2 1984.

Vichnevetsky:1984:ACM

[VS84b] R. Vichnevetsky and R. S. Stepleman, editors. Advances in computer methods for partial differential equations V: proceedings of the Fifth IMACS International Symposium on Computer Methods for Partial Differential Equations held at Lehigh University — Bethlehem, Pennsylvania, USA., June 19– 21, 1984. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1984. ISBN ???? LCCN ????

Viswanath:1998:CNR

[VT98] D. Viswanath and L. N. Trefethen. Condition numbers of random triangular matrices. SIAM Journal on Matrix Analysis and Applications, 19(2): 564–581, April 1998. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31286.

VanDeun:2011:RIC

[VT11] Joris Van Deun and Lloyd N.

Trefethen. A robust implementation of the Carathéodory– Fejér method for rational approximation. BIT Numerical Mathematics, 51(4):1039–1050, December 2011. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/content/
ag2514840142707r/;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=51&issue=4&spage=
1039.

Werner:1983:CAC

[W^{+}83] H. (Helmut) Werner et al., editors. Computational aspects of complex analysis: proceedings of the NATO Advanced Study Institute held at Braunlage, Harz, Germany, July 26– August 6, 1982, volume 102 of NATO advanced study institutes series. Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1983. ISBN 90-277-1571-

8. LCCN QA300 .N29 1982.

Wright:2015:ECP

[WJMT15] Grady B. Wright, Mohsin Javed, Hadrien Montanelli, and Lloyd N. Trefethen. Extension of Chebfun to periodic functions. SIAM Journal on Scientific Computing, 37(5):C554–C573, ???? 2015. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Weideman:1988:ESO

[WT88] J. A. C. Weideman and Lloyd N. Trefethen. The eigenvalues of second-order spectral differentiation matrices. SIAM Journal on Numerical Analysis, 25(6): 1279–1298, December 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wegert:1994:BNP

[WT94] Elias Wegert and Lloyd N. Trefethen. From the Buffon needle problem to the Kreiss matrix theorem. Amer. Math. Monthly, 101(2):132–139, 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wright:2001:CLC

[WT01a] Thomas G. Wright and Lloyd N. Trefethen. Computing Lyapunov constants for random recurrences with smooth coefficients. Journal of Computational and Applied Mathematics, 132(2):331–340, July 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042700004374.

Wright:2001:LSC

[WT01b] Thomas G. Wright and Lloyd N. Trefethen. Large-scale computation of pseudospectra using ARPACK and eigs. SIAM Journal on Scientific Computing, 23(2):591–605, March 2001. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37322. Copper

Mountain Conference (2000).

Wright:2002:PRM

[WT02] Thomas G. Wright and Lloyd N.

Trefethen. Pseudospectra of rectangular matrices. IMA J. Numer. Anal., 22(4):501– 519, October 2002. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

URL http://www3.oup.co.uk/
imanum/hdb/Volume_22/Issue_
04/220501.sgm.abs.html;
http://www3.oup.co.uk/imanum/
hdb/Volume_22/Issue_04/pdf/
220501.pdf.

Weideman:2006:OTC

[WT06] J. A. C. Weideman and L. N.

Trefethen. Optimizing Talbot’s contours for the inversion of the Laplace transform. SIAM Journal on Numerical Analysis, 44(6):2342–2362, January 2006. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Weideman:2007:KPF

[WT07a] J. A. C. Weideman and L. N. Trefethen. The kink phenomenon in Fejér and Clenshaw–Curtis quadrature. Num. Math, 107(4):707–727, October 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weideman:2007:PHC

[WT07b] J. A. C. Weideman and L. N. Trefethen. Parabolic and hyperbolic contours for computing the Bromwich integral. Mathematics of Computation, 76 (259):1341–1356, July 2007. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

                                        /www.ams.org/mcom/2007-76-
                                        259/S0025-5718-07-01945-X/
                                        home.html;   http://www.ams.
                                        org/mcom/2007-76-259/S0025-
                                        5718-07-01945-X/S0025-5718-
                                        07-01945-X.dvi;   http://
                                        www.ams.org/mcom/2007-76-
                                        259/S0025-5718-07-01945-X/
                                        S0025-5718-07-01945-X.pdf;
                                         http://www.ams.org/mcom/
                                        2007-76-259/S0025-5718-07-
                                        01945-X/S0025-5718-07-01945-
                                        X.ps.

Webb:2012:SBI

[WTG12] Marcus Webb, Lloyd N. Trefethen, and Pedro Gonnet. Stability of barycentric interpolation formulas for extrapolation. SIAM Journal on Scientific Computing, 34(6):A3009– A3015, 2012. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Xue:2024:CTD

[XWT24] Yidan Xue, Sarah L. Waters, and Lloyd N. Trefethen. Computation of two-dimensional Stokes flows via lightning and AAA rational approximation. SIAM Journal on Scientific Computing, 46(2):A1214–A1234, 2024. CODEN SJOCE3. ISSN 1064-8275,1095-7197. URL https://

doi.org/10.1137/23M1576876.